### v2

# Detecting Adversarial Data via Perturbation Forgery

Abstract
As a defense strategy against adversarial attacks, adversarial detection aims to identify and filter out adversarial data from the data flow based on discrepancies in distribution and noise patterns between natural and adversarial data. Although previous detection methods achieve high performance in detecting gradient-based adversarial attacks, new attacks based on generative models with imbalanced and anisotropic noise patterns evade detection. Even worse, existing techniques either necessitate access to attack data before deploying a defense or incur a significant time cost for inference, rendering them impractical for defending against newly emerging attacks that are unseen by defenders. In this paper, we explore the proximity relationship between adversarial noise distributions and demonstrate the existence of an open covering for them. By learning to distinguish this open covering from the distribution of natural data, we can develop a detector with strong generalization capabilities against all types of adversarial attacks. Based on this insight, we heuristically propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production, to train an adversarial detector capable of detecting unseen gradient-based, generative-based, and physical adversarial attacks, while remaining agnostic to any specific models. Comprehensive experiments conducted on multiple general and facial datasets, with a wide spectrum of attacks, validate the strong generalization of our method.
Table 1: Properties of our method vs Other Adversarial Detection: We summarize four metrics for adversarial detection. “Detect Attacks” means what kind of attacks the method is designed to detect. “Attack-Agnostic” describes whether the detection can detect unseen attack types. “Model-Agnostic” describes whether the method is independent of the protected model. Finally, “Time Cost” measures the time cost of a detector during inference.
Methods Detect Attacks Attack-Agnostic Model-Agnostic Time Cost
LID (Ma et al. 2018) Gradient $\times$ $\times$ $\ast$
LiBRe (Deng et al. 2021) Gradient $\times$ $\ast$
[…]
SPAD (Wang et al. 2023a) Gradient + GAN $\ast$
Ours Gradient + GAN + Diffusion $\ast$
Introduction
Numerous studies have demonstrated the effectiveness of deep neural networks (DNNs) in various tasks (Shi et al. 2020). However, it is also well-known that DNNs are vulnerable to adversarial attacks, which generate adversarial data by adding imperceptible perturbations to natural images, potentially causing the model to make abnormal predictions. This vulnerability reduces the reliability of deep learning systems, highlighting the urgent need for defense techniques. Some approaches employ adversarial training techniques (Zhang et al. 2019), which incorporate adversarial data into the training process to inherently bolster the model’s immunity against adversarial noise (adv-noise). Other methods utilize adversarial purification (Yoon, Hwang, and Lee 2021) to cleanse the input data of adv-noise before model inference. Unfortunately, adversarial training often results in diminished classification accuracy on adversarial data and may sacrifice performance on natural data. While purification techniques may seem more reliable, the denoising process can inadvertently smooth the high-frequency texture of clean images, leading to suboptimal accuracy on both natural and adversarial images.
Another branch of adversarial defense is adversarial detection (Feinman et al. 2017; Ma et al. 2018), which aims to filter out adversarial data before they are fed into target systems. These methods rely on discerning the discrepancies between adversarial and natural distributions, thereby maintaining the integrity of target systems while achieving high performance in defending against adversarial attacks. However, existing adversarial detection approaches primarily train specialized detectors tailored for specific attacks or classifiers (Deng et al. 2021), which may not generalize well to unseen advanced attacks.
Some recent approaches (Zhang et al. 2023; Wang et al. 2023a) have attempted to enhance the generalization of adversarial detectors by leveraging diffusion models and manually designed pseudo-noise. Nevertheless, as shown in Table 1, this improvement often comes at the cost of increased inference time, and even well-trained detectors still struggle with adversarial data produced by generative models such as generative adversarial networks (GANs) and diffusion models. Specifically, generative-based adversarial attacks, such as M3D (Zhao et al. 2023) and Diff-PGD (Xue et al. 2024), are recently developed techniques designed to generate more natural adversarial data. Unlike traditional gradient-based attacks like PGD (Madry et al. 2017), generative-based attacks tend to create imbalanced and anisotropic perturbations that are intense in high-frequency and salient areas but mild in low-frequency and background areas. These perturbations circumvent all current adversarial detection methods, posing a new challenge to the field. Therefore, there is an urgent need for a low-time-cost detection method with strong generalization performance against both unseen gradient-based attacks and unseen generative-based attacks.
In this paper, we investigate the association between the distributions of different adv-noises. As shown in Figure 1, by modeling noise as multivariant Gaussian distributions, we define the proximity relationship of the noise distributions in a metric space and establish a perturbation method to construct the proximal distributions. Based on these assumptions and definitions, we demonstrate that all adv-noise distributions are proximal using the Wasserstein distance, and we deduce a corollary that an open covering of the adv-noise distributions exists. We then propose the core idea of this paper: by learning to distinguish the open covering of adv-noise distributions from the natural distribution, we obtain a detector with strong generalization performance against all types of unseen attacks.
Figure 1: Illustration of the proximity of adv-noise distributions and perturbed noise distributions. Left: all the distributions of adv-noise and the perturbed distributions are in a $\epsilon$-ball centered on a given adv-noise distribution. Right: by continuously perturbing the commonly used adv-noise distribution, we obtain an open covering of distributions of adversarial noise.
Based on this idea, we propose Perturbation Forgery, a pseudo-adversarial data generation method designed to train a detector with strong generalization capabilities for detecting all types of adversarial attacks. As shown in Figure 2, Perturbation Forgery comprises noise distribution perturbation, sparse mask generation, and pseudo-adversarial data (pseudo-adv-data) production. By continuously perturbing the noise distribution of a commonly used attack, such as FGSM (Goodfellow, Shlens, and Szegedy 2014), we obtain a nearly complete family of distributions that form the aforementioned open covering. Simultaneously, sparse masks are generated using attention maps and saliency maps of the natural data, converting the sampled global noises into local forms to help the detector focus on local noise patterns so as to detect generative-based adv-noise. Sampling noises from the perturbed noise distributions, we sparsify these noises with the sparse masks and inject them into natural data. This process converts half of the natural data into pseudo-adv-data, which are then combined with the remaining uncontaminated data to train the binary adversarial detector. To the best of our knowledge, we are the first to empower a detector with the robust ability to detect all types of adversarial attacks, including gradient-based, GAN-based, diffusion-based, and physical attacks, with minor time cost.
Our contributions are summarized in four thrusts:
$\bullet$ By modeling adversarial noise as Gaussian distributions, we investigate the association between different adversarial noise distributions and demonstrate their proximity in a metric space.
$\bullet$ Based on our assumptions and definitions, we deduce the existence of an open covering for adv-noise distributions and propose training a detector to distinguish this open covering from the natural data distribution, thereby achieving strong generalization.
$\bullet$ Building on our deductions, we propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production. This method is designed to train a binary detector with strong generalization performance against all types of unseen attacks, including gradient-based, GAN-based, diffusion-based, and physical adversarial data, with minor time cost.
$\bullet$ We conduct evaluations on multiple general and facial datasets across a wide spectrum of gradient-based, generative-based, and physical attacks, demonstrating the consistently strong performance of our method in adversarial detection tasks.
Related Work
Adversarial Attack
Gradient-based adversarial attack
Adversarial attacks (Goodfellow, Shlens, and Szegedy 2014; Madry et al. 2017; Kurakin, Goodfellow, and Bengio 2016) attempt to fool a classifier into changing prediction results by injecting subtle perturbations to the input data while maintaining imperceptibility from human eyes. In an effort to improve transferability under the black-box setting, DIM (Wu et al. 2021), SIM (Lin et al. 2019), and VIM (Wang and He 2021) are recent attacks and achieve a high attack rate against several defense techniques.
Generative-based Adversarial Attacks
Recently, another branch of attack uses generative models to make adversarial data with a more natural style and imbalanced noise compared to gradient-based attacks. Among them, CDA (Naseer et al. 2019), TTP (Naseer et al. 2021), and M3D (Zhao et al. 2023) utilize GAN to craft adversarial data and achieve higher efficiency when attacking large-scale datasets. Diff-Attack (Chen et al. 2023) and Diff-PGD (Xue et al. 2024) employ the diffusion model as a strong prior to better ensure the realism of generated data. Designed for perturbing facial recognition systems, AdvMakeup (Yin et al. 2021) and AMT-GAN (Hu et al. 2022) are presented to generate imperceptible adversarial images of the face.
In this paper, a wide spectrum of gradient-based generative-based and adversarial attacks on ImageNet and facial datasets is used to for evaluation.
Adversarial Detection
As one of the technical solutions to ensure the safety of DNNs, there has been a significant amount of exploration in the field of adversarial detection (Ma et al. 2018). LiBRe (Deng et al. 2021) leverages Bayesian neural networks to endow pre-trained DNNs with the ability to defend against unseen adversarial attacks, SPAD (Wang et al. 2023a) leverages a data augmentation method to detect adversarial face images from unseen attack, while EPSAD (Zhang et al. 2023) incorporates a diffusion model to detect adversarial data with minimal perturbation magnitude, albeit at the cost of inference time.
While current methods have achieved remarkable improvements in adversarial detection, generative-based adversarial attacks can easily circumvent them, posing a new challenge. To address the problems above, our method is proposed to detect all types of unseen attacks, including gradient-based, generative-based, and physical attacks.
Proximity of Adversarial Noise Distribution
In this section, we provide an overview of preliminary concepts related to adversarial data generation and explore the relationship between adversarial noise distributions. We define the proximity of noise distributions and examine the imperceptibility of adversarial noise. Following this, we present the core idea of this paper: by learning to distinguish the open covering of adversarial noise distributions from the natural distribution, we develop a detector with strong generalization capabilities against all types of unseen attacks.
Adversarial data generation
Given a DNN $f$ that works on dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}^{n}_{i=1}$ with $(\mathbf{x}_{i}\in\mathcal{X},y_{i}\in\mathcal{C})$ being (sample, ground-truth) pair, an adversarial example $\hat{\mathbf{x}}$ regarding sample $\mathbf{x}$ with perturbation magnitude $\epsilon$ is calculated as:
$\hat{\mathbf{x}}=\mathop{\mathrm{argmax}}_{\hat{\mathbf{x}}\in\mathbb{B}(\mathbf{x},\epsilon)}\mathcal{L}(f(\hat{\mathbf{x}}),y),$ (1)
where $\mathcal{L}$ is some loss function and $\mathcal{B}(\mathbf{x},\epsilon)=\{\mathbf{x^{\prime}}|\ell^{p}(\mathbf{x},\mathbf{x^{\prime}})<\epsilon\}$ restricts the perturbation under $\ell^{p}$ norm. Following (Goodfellow, Shlens, and Szegedy 2014), we simply denote adversarial sample as $\hat{\mathbf{x}}=\mathbf{x}+\boldsymbol{\eta}$.
Proximity of noise distribution
Due to the subtle and imperceptible nature of adv-noise, adversarial data and natural data appear visually similar. Thus, we assume that this similarity also exists between the noises of different adversarial attacks. To investigate the similarity of these noises, we provide the following definition, which depicts a close relationship between two independent noise distributions:
Definition 1 (Proximal Noise Distribution).
Let $P$ and $Q$ be the distribution functions of two independent noise samples $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$. We call them proximal noise distributions under $\varepsilon$ if we have a metric $d(\cdot,\cdot)$ and $\exists\ \varepsilon>0$, s.t.
$d(P,Q)<\varepsilon.$ (2)
A common assumption for instance feature distribution is to model it as a Gaussian distribution (Du et al. 2022). Obtained by subtracting from clean data and adversarial data, adv-noises naturally form a Gaussian distribution. Therefore, we assume that noise r.v $\boldsymbol{\eta}$ satisfied truncated Gaussian distribution $\mathcal{N}(\mu,\sigma^{2}\mathbf{I}_{d})$ with it’s truncated interval is $[-\epsilon,\epsilon]^{d}$. Built upon Definition 1, we derive the following theorem about the association between the distributions of different adv-noises:
Theorem 1.
Let $\mathcal{P}_{a}$ be the distribution set composed of all the adversarial noise distributions. Given independent noise distributions $Q_{i},i\in N^{+}$. For $\forall\ i\neq j$, $Q_{i}$ and $Q_{j}$ are proximal noise distributions if the following conditions are met.
1) if $Q_{i},Q_{j}\in\mathcal{P}_{a}$.
[…]
Theorem 1 tells us: 1) All adversarial noise distributions are proximal noise distributions. 2) If the parameters of two noise distributions are sufficiently close to those of a given distribution, we can regard these two distributions as proximal noise distributions.
Figure 2: Illustration of Perturbation Forgery. Before training, we estimate the noise distribution of a commonly used attack and continuously perturb it in each batch. After that, we sample noises from the perturbed distribution and convert them into local noise using sparse masks. Finally, pseudo-adversarial data is generated by adding these local noises to the natural samples.
We can further deduce several corollaries of the theorem. 1) All noise distributions proximal to the known adversarial distribution $p(\boldsymbol{\eta})$ form an open $\varepsilon$-ball centered on $p(\boldsymbol{\eta})$, denoting as $\mathcal{P}_{\varepsilon}=\{p(\boldsymbol{\eta}_{i})|d(p(\boldsymbol{\eta}_{i}),p(\boldsymbol{\eta}))<\varepsilon)\}$, and also form a metric space $(\mathcal{P}_{\varepsilon},d)$. 2) The adv-noise distribution set is also located in this ball: $\mathcal{P}_{a}\subset\mathcal{P}_{\varepsilon}$. 3) Based on the properties of the metric space $(\mathcal{P}_{\varepsilon},d)$, for all subset $\mathcal{P}_{i}\subset\mathcal{P}_{\varepsilon}$, there exists an open covering of the subset: $\mathcal{P}_{i}\subset\{G_{c}\}_{c\in J}$.
Based on the corollaries, we deduce the core idea of this paper: Let $p(\hat{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})=\mathcal{N}(\mu+\epsilon_{\mu i},(\sigma+\epsilon_{\sigma i})^{2}\mathbf{I}_{d}),i=1,2,\dots,n$ denote generated proximal noise distributions of the given adv-noise distribution $p(\boldsymbol{\eta})=\mathcal{N}(\mu,\sigma^{2}\mathbf{I}_{d})$ where $\epsilon_{\mu i},\epsilon_{\sigma i}$ are randomly sampled from uniform distributions and restricted by $\epsilon_{\mu i},\epsilon_{\sigma i}<\varepsilon$. As shown in Figure 1, when $n\rightarrow\infty$, we can always obtain a distribution set $\mathcal{P}_{c}=\{p(\hat{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})\}_{i=1}^{\infty}$ that forms an open covering of the adversarial distribution set $\mathcal{P}_{a}$. Trained with the open covering, the detector can distinguish the distributions of natural data and adversarial data, thereby having the functionality to detect unseen adversarial attacks.
For proofs of theorem and corollaries, please refer to the supplementary materials.
Perturbation Forgery
Based on the theory support in the last section, we heuristically propose Perturbation Forgery to train a detector with strong generalization against all types of unseen attacks.
As illustrated in Figure 2, Perturbation Forgery encompasses three elements: distribution perturbation, sparse mask generation, and pseudo-adversarial data production. At first, the noise distribution of a commonly used attack is continuously perturbed using Noise Distribution Perturbation into numerous noise distributions that remain proximal to the original, thereby forming the open covering. After that, we sample noises from the perturbed distribution and convert them into sparse local noise using Sparse Mask Generation. Subsequently, we convert half of the natural data into pseudo-adversarial data using Pseudo-Adversarial Data Generation and train a binary detector on the these data along with the remaining natural data.
Noise Distribution Perturbation
Given the training dataset $\{(\mathbf{x}_{k},y_{k})\}^{n}_{k=1}$ composed by natural data, we add adv-noise on it using a commonly used attack (such as FGSM) and generate adversarial noise features $\mathcal{D}_{\mathrm{noise}}:=\{\mathbf{z}_{k}=h(\boldsymbol{\eta}_{k})\}_{k=1}^{n}$ by
$\boldsymbol{\eta}_{k}=\operatorname{Attack}(\mathbf{x})-\mathbf{x},\ k\in[1,n],$ (3)
where $h:\mathcal{X}\rightarrow\mathbb{R}^{d}$ is the flattening operation that converts multidimensional data into vectors.
Assuming the noise features form a multivariate Gaussian distribution $p(\mathbf{z})=\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ (Du et al. 2022), we estimate the parameters mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$ of $\mathcal{D}_{\mathrm{noise}}$ by
$\widehat{\boldsymbol{\mu}}=\frac{1}{N}\sum_{k}\mathbf{z}_{k},\quad\widehat{\boldsymbol{\Sigma}}=\frac{1}{N}\sum_{k}\left(\mathbf{z}_{k}-\widehat{\boldsymbol{\mu}}\right)\left(\mathbf{z}_{k}-\widehat{\boldsymbol{\mu}}\right)^{\top}.$ (4)
For the $i^{\mathrm{th}}$ batch of training, we perturb the estimated distribution to a proximal noise distribution by
$\widehat{\boldsymbol{\mu}}_{i}=\widehat{\boldsymbol{\mu}}+\alpha_{i}\cdot\boldsymbol{m}_{i},\quad\widehat{\boldsymbol{\Sigma}}_{i}=\widehat{\boldsymbol{\Sigma}}+\beta_{i}\cdot\boldsymbol{V}_{i},$ (5)
where $\boldsymbol{m}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ and $\boldsymbol{V}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d\times d})$ are vector and matrix of perturbation. $\alpha_{i}\sim\boldsymbol{U}(-\epsilon_{\mu},\epsilon_{\mu})$ and $\beta_{i}\sim\boldsymbol{U}(0,\epsilon_{\sigma})$ are scale parameters sampled from uniform distributions. After $n$ batches of training, we have a perturbed distribution set $\{\mathcal{N}(\widehat{\boldsymbol{\mu}}_{i},\widehat{\boldsymbol{\Sigma}}_{i})\}$, thereby forming the approximate open covering $\mathcal{P}_{c}$. Trained with $\mathcal{P}_{c}$, the detector can distinguish the distributions of natural and adversarial data.
Sparse Mask Generation
Unlike gradient-based adversarial attacks, generative-based attacks tend to generate imbalanced and anisotropy perturbations, showing much intensity in high-frequency and salient areas and mild in low-frequency and background areas. Besides, in the physical situation, adversarial attacks also tend to add patches to local areas. Noticing this, we propose sparse masks to perturbation forgery to assist the detector in detecting local adv-noises.
Previous research has demonstrated that the spectra of real and fake images differ significantly in high-frequency regions (Luo et al. 2021). Additionally, generative-based adversarial data manipulate characteristics, such as abnormal color aberrations, in these high-frequency regions (Wang et al. 2023a). To augment the detection performance on generative-based adversarial data, we need to sparse the pseudo-noise sampled from perturbed noise distribution and highlight the noise in high-frequency regions to mimic the style of generative-based attacks.
Algorithm 1 Training Procedure with Perturbation Forgery
Input: Training dataset, initial attack $\operatorname{Attack}$, saliency detection model $f_{s}$, and GradCAM model $f_{c}$.
Output: A detector with strong generalization capabilities against all types of unseen attacks.
1:  Generate adversarial noise using the initial attack $\operatorname{Attack}$ in Eq. 3.
2:  Estimate a multivariate Gaussian distribution $p(\mathbf{z})=\mathcal{N}(\widehat{\boldsymbol{\mu}},\widehat{\boldsymbol{\Sigma}})$ of adv-noise using Eq. 4.
3:  for batch $i$ in $1,\cdots,n$ do
4:     Generate a perturbed noise distribution $p(\mathbf{z}_{i})=\mathcal{N}(\widehat{\boldsymbol{\mu}}_{i},\widehat{\boldsymbol{\Sigma}}_{i})$ using Eq. 5.
5:     for sample $\mathbf{x}_{i,k}$ half of batch $i$ do
6:        Generate sparse masks $\operatorname{Mask}(\mathbf{x}_{i,k})$ using Eq. 8.
7:        Sample pseudo-adversarial noise $\widehat{\boldsymbol{\eta}}_{i,k}$ using Eq. 10.
8:        Generate pseudo-adversarial data $\widehat{\mathbf{x}}_{i,k}$ using Eq. 12.
9:     end for
10:     Train the detector on batch $i$.
11:  end for
Given this, we first mask the high-frequency area of a natural sample using saliency detection (Qin et al. 2019) and GradCAM (Selvaraju et al. 2017):
$\operatorname{Mask}_{1}(\mathbf{x})=\operatorname{Map}(f_{s}(\mathbf{x});\gamma_{s})\cup\operatorname{Map}(f_{c}(\mathbf{x});\gamma_{c}),$ (6)
where $f_{s}$ and $f_{c}$ are the saliency detection model and GradCAM model respectively. $\operatorname{Map}(\mathbf{x};\gamma)$ is a binary indicator function parameterized by $\gamma$:
$\operatorname{Map}(\mathbf{x};\gamma)=\operatorname{Clip}_{[0,1]}(\mathbf{x}-\gamma\mathbf{I}_{\mathcal{X}}).$ (7)
The noise that occurs in the high-frequency region tends to occupy only a small part of the region rather than spreading to the entire high-frequency spectrum. Therefore, we select the high-frequency point using the Sobel operator (Duda and Hart 1973) to sparsify the mask of sample $\mathbf{x}_{i,k}$:
$\operatorname{Mask}_{2}(\mathbf{x})=\operatorname{Map}(\operatorname{Sobel}(\mathbf{x});\gamma_{l}),$ (8)
$\operatorname{Mask}(\mathbf{x}_{i,k})=\operatorname{Mask}_{1}(\mathbf{x}_{i,k})\cap\operatorname{Mask}_{2}(\mathbf{x}_{i,k}).$ (9)
Pseudo-Adversarial Data Production
For the $i^{\mathrm{th}}$ batch of training, we sample pseudo-adversarial noise from the perturbed distribution by:
$\widehat{\boldsymbol{\eta}}_{i,k}=h^{-1}(\mathbf{z}_{i,k}),$ (10)
$\mathbf{z}_{i,k}\in\{\mathbf{z}_{k}|\mathbf{z}_{k}\sim\mathcal{N}(\widehat{\boldsymbol{\mu}}_{i},\widehat{\boldsymbol{\Sigma}}_{i})\ \mathrm{and}\ p(\mathbf{z}_{k})<\gamma_{p}\}$ (11)
where $h^{-1}:\mathbb{R}^{d}\rightarrow\mathcal{X}$ is an inverse function of $h$ to unflatten noise features, and $\gamma_{p}$ is a threshold that excludes data with a low likelihood. Note that $\boldsymbol{\eta}_{i}$ is a global noise and needs to be localized by a sparse mask. For natural data $\mathbf{x}_{i,k}$ in the $i^{\mathrm{th}}$ batch, the corresponding pseudo-adversarial data $\widehat{\boldsymbol{x}}_{i,k}$ is produced by adding the masked local noise to it:
$\widehat{\mathbf{x}}_{i,k}=\mathbf{x}_{i,k}+\widehat{\boldsymbol{\eta}}_{i,k}\otimes\operatorname{Mask}(\mathbf{\mathbf{x}_{i,k}}).$ (12)
where $\otimes$ is Hadamard product.
For each batch, we convert half of the natural data to pseudo-adversarial data by Perturbation Forgery and train a binary detector using the Cross-Entropy classification loss (Zhang and Sabuncu 2018). The overall training procedure with Perturbation Forgery is presented in Algorithm 1.
Experiments
To validate the detection performance of our approach against various adversarial attack methods, we conduct extensive experiments on multiple general and facial datasets.
Experiment Settings
Datasets and Network Architectures
The datasets for evaluation include 1) general dataset ImageNet100 (Deng et al. 2009), and 2) face datasets: Makeup (Gu et al. 2019), CelebA-HQ (Karras et al. 2017), and LFW (Gary et al. 2007). Gradient-based attacks and generative-based attacks are launched only on ImageNet100. For face attacks, we conduct Adv-Makeup on Makeup, AMT-GAN on CelebA-HQ, and physical attacks on LFW.
Adversarial Attacks
We adopt 9 gradient-based attacks, 5 generative-based attacks, and 6 face attacks to evaluate our method. For gradient-based attacks, BIM (Kurakin, Goodfellow, and Bengio 2016), PGD (Madry et al. 2017), RFGSM (Tramèr et al. 2018), MIM (Dong et al. 2018), DIM (Wu et al. 2021), NIM and SNIM (Lin et al. 2019), VNIM (Wang and He 2021), and Auto-Attack (Croce and Hein 2020) are used as the unseen attacks. For generative-based attacks, we adopt 3 GAN-based attacks including CDA (Naseer et al. 2019), TTP (Naseer et al. 2021), and M3D (Zhao et al. 2023), and 2 Diffusion-based attacks including Diff-Attack (Chen et al. 2023) and Diff-PGD (Xue et al. 2024) to craft adv-samples. For face attacks, 1 gradient-based attack TIPIM (Yang et al. 2021), 2 GAN-based attacks including AdvMakeup (Yin et al. 2021) and AMT-GAN (Hu et al. 2022), and 3 physical attacks including Adv-Sticker (Tong et al. 2021), Adv-Glasses (Sharif et al. 2016), and Adv-Mask (Tong et al. 2021) produce adversarial face samples. For each of the above attack methods, adversarial data is generated with the perturbation magnitude of $\epsilon=4/255$ (except for the unrestricted attacks) following (Zhang et al. 2023).
Table 2: Comparison of AUC scores (%) of detecting gradient-based attacks on ImageNet100.
Detector BIM PGD RFGSM DIM MIM NIM VNIM SNIM AA
LID (Ma et al. 2018) 0.9782 0.9750 0.9637 0.8942 0.9146 0.8977 0.8865 0.8752 0.9124
LiBRe (Deng et al. 2021) 0.9259 0.9548 0.9769 0.9243 0.8725 0.9013 0.8862 0.8694 0.8653
SPAD (Wang et al. 2023a) 0.9846 0.9851 0.9833 0.9815 0.9820 0.9823 0.9891 0.9877 0.9890
EPSAD (Zhang et al. 2023) 0.9998 0.9989 0.9992 0.9923 0.9918 0.9972 0.9946 0.9908 0.9998
ours 0.9911 0.9912 0.9911 0.9863 0.9931 0.9934 0.9878 0.9914 0.9941
Table 3: Comparison of AUC scores of detecting generative-based attacks on ImageNet100.
Detector CDA TTP M3D Diff-Attack Diff-PGD
CNN-Detection (Wang et al. 2020) 0.7051 0.6743 0.6917 0.3963 0.526
LGrad (Tan et al. 2023) 0.6144 0.6077 0.6068 0.5586 0.5835
Universal-Detector (Ojha, Li, and Lee 2023) 0.7945 0.8170 0.8312 0.9202 0.5531
DIRE (Wang et al. 2023b) 0.8976 0.9097 0.9129 0.9097 0.9143
EPSAD 0.9674 0.6997 0.7265 0.4700 0.1507
SPAD 0.9385 0.9064 0.8984 0.8862 0.8879
ours 0.9878 0.9678 0.9364 0.9223 0.9223
Baselines
We compare our method with state-of-the-art detection methods across three research fields, including adversarial detection, synthetic image detection, and face forgery detection, evaluating the detection of various types of adversarial attacks. For detecting gradient-based attacks, we compared our method with adversarial detection methods including LID (Ma et al. 2018), LiBRe (Deng et al. 2021), SPAD (Wang et al. 2023a), and EPSAD (Zhang et al. 2023). Since existing adversarial detection methods do not address generative-based adversarial data, we compared our method with synthetic image detection methods including CNN-Detection(Wang et al. 2020), LGrad(Tan et al. 2023), Universal-Detector(Ojha, Li, and Lee 2023), and DIRE(Wang et al. 2023b) for detecting generative-based attacks. For detecting GAN-based face attacks, we compared with face forgery detection methods including Luo et al. (Luo et al. 2021), He et al. (He et al. 2021), and ODIN (Liang, Li, and Srikant 2017). We conducted experiments using the pre-trained models and authors’ implementation code of the above methods.
Evaluation Metrics
Following previous adversarial detection works (Zhang et al. 2023), we employ the Area Under the Receiver Operating Characteristic Curve (AUC) as the primary evaluation metric to assess the performance of our classification models. The AUC is a widely recognized metric in the field of machine learning and is particularly useful for evaluating binary classification problems. For fair comparisons, the number of natural data and adversarial data in the experiments remained consistent.
Implementation details
XceptionNet (Chollet 2017) is selected as our backbone model and is trained on all the above datasets. For adv-noise distribution estimation, we utilize Torchattacks (Kim 2020) to generate 1000 noise images for each dataset. The uniform distribution parameters are set as $\epsilon_{\mu}=3$ and $\epsilon_{\sigma}=0.005$. The threshold of the sample control is set as $\gamma_{s}=0.05$. The thresholds of the binary indicator function are set as $\gamma_{s}=0$, $\gamma_{c}=125$ and $\gamma_{l}=100$. FGSM (Goodfellow, Shlens, and Szegedy 2014) is used as the initial attack for Perturbation Forgery. Training epochs are set to 10 and convergence is witnessed.
Detecting Gradient-based Attacks
We begin by comparing our method with state-of-the-art techniques in detecting gradient-based attacks on ImageNet100. All the attacks are seen to LID but unseen to other baselines and our method. For each attack, 1,000 adversarial images are generated in the white-box setting to test the detection models. All baseline methods are sourced from existing adversarial detection research. As reported in Table 2, although EPSAD performs well, our method demonstrates effective detection (achieving an AUC score of 0.9931 against MIM and 0.9914 against SNIM) across a variety of attack methods. These results validate our assumptions about the proximity of the noise distribution. By continuously perturbing the estimated noise distribution of a given attack (FGSM), we obtain an open covering noise of distributions from all the attacks. Given one seen attack, we are able to detect all the unseen attacks. The consistent detection performance verifies the effectiveness of our core idea of training a detector to distinguish the open covering of adv-noise distributions from natural ones and demonstrates the strong generalization capability of our method against gradient-based attacks. Additional results on CIFAR10 are provided in the supplementary materials.
Table 4: Comparison of AUC scores for detecting various adversarial attacks on the face.
Detector TIPIM Adv-Sticker Adv-Glasses Adv-Mask Detector Adv-Makeup AMT-GAN
LID 0.6974 0.5103 0.6472 0.5319 He et al. (He et al. 2021) 0.5252 0.8823
LiBRe 0.5345 0.9739 0.8140 0.9993 Luo et al. (Luo et al. 2021) 0.6178 0.6511
EPSAD 0.9647 0.9462 0.9260 0.9578 ODIN(Liang, Li, and Srikant 2017) 0.6325 0.6970
SPAD 0.9256 0.9975 0.9244 0.9863 SPAD 0.9657 0.8969
ours 0.9885 0.9999 0.9999 0.9999 ours 0.9762 0.9216
Detecting generative-based Attacks
To investigate the advantage of our method in detecting generative-based attacks, we compare with state-of-the-art methods of synthetic image detection and adversarial detection methods on ImageNet100. As shown in Table 3, our method substantially surpasses other methods, exhibiting highly effective detection capabilities. For adversarial detection baselines, due to the noise produced by generative models tending to be imbalanced compared to perturbations of traditional gradient-based attacks, these methods struggle to conduct effective detection. In contrast, with the assistance of sparse masks, our detector can perceive the local noises brought by GAN and diffusion models. As adversarial data is crafted from natural data, most synthetic image detection methods lose their effectiveness except DIRE(Wang et al. 2023b). This is probably because DIRE leverages the reconstruction error of images generated by diffusion models, thus maintaining a certain level of effectiveness.
Table 5: Ablation study on masks.
Ablation CDA TTP M3D Diff-Attack Diff-PGD
w/o $f_{s}$ 0.9755 0.9585 0.9320 0.9197 0.9135
w/o $f_{c}$ 0.9864 0.9510 0.9262 0.9031 0.9062
w/o $\operatorname{Sobel}$ 0.9868 0.7235 0.6984 0.5231 0.5547
Table 6: Time cost of inference (second) on ImageNet100
Detector LID LiBRe EPSAD SPAD ours
Time (second) 1.80 2.56 396.81 4.56 4.85
Detecting Face Adversarial Examples
Adversarial attack on the face recognition model is an important branch of adversarial example research. Therefore, the detection ability of adversarial face examples is an important indicator of the performance of the detector. We perform various adversarial face attacks and compare our approach with state-of-the-art methods from adversarial face detection and face forgery detection. As shown in Table 4, our method shows excellent performance against a variety of adversarial attacks targeting faces and surpasses many existing adversarial face detection techniques. It further indicates the ability of the sparse mask to capture the essential characteristics of face adversarial examples and the strong generalization capability of our method.
Ablation Study and Analysis
To further evaluate our method, we conduct experiments to show the effect of the sparse mask and compare the time cost to baseline methods. We also analyze the functionality of Perturbation Forgery through visualizations. The impact of hyper-parameters is reported in the supplementary materials.
Impact of each mask
As argued, the sparse mask is proposed to assist the detector in identifying local adversarial noises present in generative-based adversarial data. As shown in Table 5, the absence of the saliency detection model $f_{s}$ or the GradCAM model $f_{c}$ has little effect on the overall system. This is because the regions masked by $f_{s}$ and $f_{c}$ are somewhat complementary. However, without the Sobel operation to make the mask sparse, the detector fails entirely in detecting attacks, except for CDA, because the noise in generative-based adversarial data tends to be imbalanced and localized, whereas the noise in CDA is more global. Without the sparsification provided by the Sobel operation, the generated pseudo-noise becomes too dense, preventing the detector from effectively focusing on local noise.
Time Cost of Inference
We conducted experiments to calculate the inference time cost for 100 samples on ImageNet100. As shown in Table 6, the time cost of our model is slightly higher than that of LID, LiBRe, and SPAD, and it only takes 0.0485 seconds to process one image. For actual use, a minor increase in time is perfectly acceptable in exchange for a significant increase in detection performance.
Figure 3: 2D T-SNE visualizations. (a) ImageNet100 flattened noises of adversarial data and Perturbation Forgery. (b) ImageNet100 features extracted by the model trained with Perturbation Forgery.
Functionality of Perturbation Forgery
To validate the functionality of our method, we extract features and flattened noises from adversarial data and Perturbation Forgery and visualize them using 2D t-SNE projection. All adversarial data come from unseen attacks. As shown in Figure 3(a), by continuously perturbing the adversarial noise distribution of a commonly used attack, Perturbation Forgery forms an open covering of noises from unseen attacks. Individual samples float outside of the collective because they are outliers from the noise distribution. As shown in Figure 3(b), by learning to separate natural data from those generated by Perturbation Forgery, the trained detector can effectively distinguish between natural and adversarial data.
Conclusion
In this paper, we explore and define the proximity relationship between adversarial noise distributions by modeling these distributions as multivariate Gaussian. Based on our assumptions and definitions, we establish a perturbation method to craft proximal distributions from a given distribution and demonstrate the existence of an open covering of adversarial noise distributions. After that, we propose the core idea of obtaining a detector with strong generalization abilities against all types of adversarial attacks by learning to distinguish the open covering from the distribution of natural data. Based on this insight, we heuristically propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production, to train an adversarial detector capable of detecting unseen gradient-based, generative-based, and physical adversarial attacks, agnostic to any specific models. The major limitation of our method is that finding the optimal scale parameters for distribution perturbation is challenging. We must rely on continuous experimentation to determine these parameters. However, even if these parameters are not optimal, our method still maintains consistent detection performance across various attacks on multiple general and facial datasets, with a satisfactory inference time cost.
Appendix
Appendix A Proofs
Theorem Proofs
Proof of Theorem 1.
Proof.
Denoting the distribution of natural sample $\mathbf{x}$ as $P$, we use the 1-Wasserstein distance as the metric $d(\cdot,\cdot)$, to prove the theorem.
1) We denote the independent adv-noise distribution of $\eta_{i}$ as $Q_{i}\in\mathcal{P}_{a}$ where $\eta_{i}=\hat{\mathbf{x}}_{i}-\mathbf{x}$, and a zero-distribution as $P$ where its mean and variance are zero vectors. Then the 1-Wasserstein distance between $\eta_{i}$ and $\eta_{j}$ can be bound by its metric property and dual form:
$W_{1}\left(Q_{i},Q_{j}\right)\leq$ $W_{1}(Q_{i},P)+W_{1}(Q_{j},P)$
$\leq$ $\sup_{\|g\|_{L}\leq 1}\mathbb{E}_{\boldsymbol{\eta}_{i}\sim Q_{i}}[g(\boldsymbol{\eta}_{i})]-\mathbb{E}_{\mathbf{x}\sim P}[g(\mathbf{x})]$
$+\sup_{\|g\|_{L}\leq 1}\mathbb{E}_{\boldsymbol{\eta}_{j}\sim Q_{j}}[g(\boldsymbol{\eta}_{j})]-\mathbb{E}_{\mathbf{x}\sim P}[g(\mathbf{x})]$
$\leq$ $2\epsilon,$
where $\|g\|_{L}$ is the Lipschitz constant of function $g$ w.r.t. the norm induced by the metric here.
2) Consider two random noise $\boldsymbol{\eta}_{i}\sim\mathcal{N}_{i}(\mu_{i},\sigma_{i}^{2}\mathbf{I}_{d})$ and $\boldsymbol{\eta}_{j}\sim\mathcal{N}_{j}(\mu_{j},\sigma_{j}^{2}\mathbf{I}_{d})$. While it’s mean and variance satisfied that $\|\mu_{i}-\mu_{j}\|\leq\epsilon$ and $|\sigma_{i}-\sigma_{j}|\leq\epsilon$. then the 2-Wasserstein distance between $\eta_{1}$ and $\eta_{2}$ is
$W_{2}\left(\boldsymbol{\eta}_{i},\boldsymbol{\eta}_{j}\right)^{2}=$ $\left\|\mu_{i}-\mu_{j}\right\|^{2}+\operatorname{trace}(\sigma_{i}^{2}\mathbf{I}_{d}+\sigma_{j}^{2}\mathbf{I}_{d}$
$-2\left((\sigma_{j}^{2}\mathbf{I}_{d})^{1/2}\sigma_{i}^{2}\mathbf{I}_{d}(\sigma_{j}^{2}\mathbf{I}_{d})^{1/2}\right)^{1/2})$
$\leq$ $\epsilon_{\mu}^{2}+d\cdot\epsilon_{\sigma}^{2},$
where $(\sigma_{i}^{2}\mathbf{I}_{d})^{1/2}$ denotes the principal square root of $\sigma_{i}^{2}\mathbf{I}_{d}$. ∎
Corollary 1.
1) All noise distributions proximal to the known adversarial distribution $p(\boldsymbol{\eta})$ form an open $\varepsilon$-ball centered on $p(\boldsymbol{\eta})$, denoting as $\mathcal{P}_{\varepsilon}=\{p(\boldsymbol{\eta}_{i})|d(p(\boldsymbol{\eta}_{i}),p(\boldsymbol{\eta}))<\varepsilon)\}$, and also form a metric space $(\mathcal{P}_{\varepsilon},d)$.
2) The adv-noise distribution set is also located in this ball: $\mathcal{P}_{a}\subset\mathcal{P}_{\varepsilon}$.
3) Based on the properties of the complete metric space $(\mathcal{P}_{\varepsilon},d)$, for all subset $\mathcal{P}_{i}\subset\mathcal{P}_{\varepsilon}$, there exists a finite open covering of the subset: $\mathcal{P}_{i}\subset\{G_{c}\}_{c\in J}$.
Proof.
1) Consider 1-Wasserstein distance as metric $d$, for $\forall Q_{i},Q_{j}\in\mathcal{P}_{\varepsilon}$, we have
$W_{1}\left(Q_{i}\right)\geq 0.\quad W_{1}\left(Q_{i},Q_{j}\right)=0\ if\ and\ only\ if\ Q_{i}=Q_{j}.$ (13)
$W_{1}\left(Q_{i},Q_{j}\right)=W_{1}\left(Q_{j},Q_{i}\right).$ (14)
$W_{1}\left(Q_{i},Q_{j}\right)\leq W_{1}(Q_{i},P)+W_{1}(Q_{j},P).$ (15)
Therefore, $(\mathcal{P}_{\varepsilon},d)$ is a metric space.
2) It is obvious according to the definition of $\mathcal{P}_{\varepsilon}$.
3) For any subset $\mathcal{P}_{i}=\{p_{i1},p_{i2},\dots,p_{in}\}\subset\mathcal{P}_{\varepsilon}$, we can construct several spherical subsets $\{\mathcal{P}_{ij}\}$ with $p_{i1}$ as the center and $|\varepsilon-d(p_{i1},p(\boldsymbol{\eta}))|$ as the radius. $\{\mathcal{P}_{ij}\}$ is a open covering of $\mathcal{P}_{i}$.
Appendix B Experimental Details
Dataset
CIFAR10
CIFAR10 consists of a training set of 50,000 images and a test set of 10,000 images, each with a resolution of 32x32 pixels. These images are divided into 10 different classes. The classes represent everyday objects such as airplanes, automobiles, birds, cats, deer, dogs, frogs, horses, ships, and trucks.
ImageNet100
ImageNet-100 is a subset of ImageNet (Deng et al. 2009), comprising 100 classes selected from the original ImageNet dataset, each containing a substantial number of high-resolution images, with 1,300 images per class.
Makeup
Makeup comprises images of faces with and without makeup. These images are collected from various sources to ensure a diverse representation of facial features, makeup styles, and skin tones. There are 333 before-makeup images and 302 after-makeup images.
Table 7: Detection AUC(%) against various gradient-based attacks on CIFAR-10 ($\epsilon=4/255$).
Detector BIM PGD RFGSM DIM MIM NIM VNIM SNIM AA
SPAD (Wang et al. 2023a) 99.51 99.38 99.35 99.52 99.66 99.85 99.47 99.63 99.74
EPSAD (Zhang et al. 2023) 99.99 99.98 99.98 99.99 99.99 99.99 99.99 99.99 99.99
ours 99.65 99.65 99.68 99.47 99.80 99.80 99.76 99.64 99.88
Figure 4: The influence of parameters on the detection performance.
Figure 5: The influence of attack intensity on the detection performance.
CelebA-HQ
CelebA-HQ dataset is a high-quality version of the original CelebA dataset.CelebA-HQ consists of 30,000 high-resolution images of celebrity faces, derived from the CelebA dataset through a progressive GAN-based upsampling and quality enhancement process. Each image in CelebA-HQ is 1024x1024 pixels.
LFW
LFW contains 13,233 labeled images of 5,749 different individuals collected from the internet, each image is labeled with the name of the person pictured. LFW images vary widely in terms of lighting, facial expression, pose, and background, closely reflecting real-world conditions. The images are provided in a resolution of 250x250 pixels, and the faces are roughly aligned based on the eye coordinates.
Adversarial Attacks
Gradient-based attacks
We use Torchattacks (Kim 2020) to implement adversarial attacks. The victim models used for adversarial data generation are ResNet-50 (He et al. 2016) for ImageNet, WideReseNet-28 for CIFAR10 (Krizhevsky, Hinton et al. 2009), and ArcFace (Deng et al. 2019) for face attacks.
Generative-model-based attacks
For CDA, we use the pre-trained generator which has been trained on ImageNet with a relativistic loss against ResNet152. For TTP, we use the pre-trained-generators trained against ResNet50 for 8 target labels, they are: Grey Owl(24),Goose(99),French Bulldog(245),Hippopotamus(344),Cannon(471),Fire Engine(555),Parachute(701), Snowmobile(802). We generate 1000 adversarial examples for each label and randomly select 125 adversarial examples for each label. For M3D, its settings are consistent with TTP. For Diff-PGD, we use the global attack to craft adversarial examples, target classifier is ResNet50, the diffusion model accelerator is ddim50, the reverse step in SDEdit is 3, and the iteration number of PGD is 10, the step size is 2. For Diff-attack, the target classifier is ResNet50, DDIM sample steps are set as 20, and iterations to optimize the adversarial image are set as 30. For the methods mentioned above, we all use the authors’ implementation to craft adversarial examples.
Face attacks
For face attacks, 1 gradient-based attack TIPIM (Yang et al. 2021), two GAN-based attacks including AdvMakeup (Yin et al. 2021) and AMT-GAN (Hu et al. 2022), and 3 physical attacks including Adv-Sticker (Tong et al. 2021), Adv-Glasses (Sharif et al. 2016), and Adv-Mask (Tong et al. 2021) produce face adversarial samples by attacking the victim face model ArcFace (Deng et al. 2019).
Baselines
Adversarial detection
For LiBRe [^1] https://github.com/thudzj/ScalableBDL/tree/efficient/exps, we trained a ResNet-50 on ImageNet100 and finetuned it using the authors’ implementation, the parameters were kept consistent with theirs. For EPSAD [^2] https://github.com/ZSHsh98/EPS-AD, we use the pre-trained models provided by the authors and follow the instructions by the authors to get the detection results. Due to the time-consuming nature of computing EPS, we evaluated 1,000 images for each attack method on ImageNet-100.
Synthetic image detection
Since existing adversarial example detection methods do not address adversarial examples based on generative models, we compared our method with various synthetic image detection methods to effectively validate its effectiveness. We conducted detection experiments on adversarial examples based on generative models using the following methods: CNN-Detection(Wang et al. 2020) [^3] https://github.com/peterwang512/CNNDetection, LGrad(Tan et al. 2023) [^4] https://github.com/peterwang512/CNNDetection, Universal-Detector(Ojha, Li, and Lee 2023) [^5] https://github.com/Yuheng-Li/UniversalFakeDetect?tab=readme-ov-file#weights, and DIRE(Wang et al. 2023b) [^6] https://github.com/ZhendongWang6/DIRE. We conducted experiments using the pre-trained models and authors’ implementation code of the above methods.
Implementation Details
Experiments are implemented using 2 Nvidia Deforce RTX 3090 GPUs. For training detectors, we use an Adam optimizer with a learning rate of 2e-4, momentum of 0.9, and weight decay of 5e-6.
Data pre-processing
We resize images to $32\times 32$ for CIFAR10, $256\times 256$ for ImageNet100, Makeup and CelebA-HQ, and $112\times 112$ for LFW. For ImageNet100, after resizing, we center-crop the images to $224\times 224$. After that, all of the images are resized to $256\times 256$ followed by random horizontal flip and normalization.
Noise distribution
As related in the main paper, we flatten the generated noise to a vector and estimate its distribution. However, for a noise with a shape of $256\times 256\times 3$, the length of the corresponding flattened noise is $196,608$, making it impossible to run the algorithm on a single GPU. Therefore, we fix the size of noise to $32\times 32\times 3$, and concatenate several sampled pseudo noises to fit the larger image.
Figure 6: Natural images and corresponding pseudo-adversarial images generated by Perturbation Forgery.
Experiment on CIFAR-10
We conduct experiments to detect gradient-based attacks on CIFAR10. As reported in Table 7, our method demonstrates effective detection (achieving an AUC score over 0.99) across a variety of attack methods, validating our assumptions of the noise distribution proximity.
Impact of Parameters
We select typical attacks and present curves in Figure 4 to depict the impact of parameters on the detection performance. The parameters $\epsilon_{\mu}$ and $\epsilon_{\sigma}$ have little effect on detecting PGD, while $\gamma_{c}$ and $\gamma_{l}$ have significant effects on detecting M3D and Diff-PGD. This might be because, regardless of whether $\epsilon_{\mu}$ and $\epsilon_{\sigma}$ are large or small, the accuracy of forming the open coverage remains unaffected as long as the data is sufficient. However, too small thresholds lead to adding pseudo-noise to insensitive and low-frequency regions, while too large thresholds result in extremely sparse pseudo-noise that is far from the adversarial noise distribution.
To explore the impact of perturbation intensity of attacks, we select typical attacks and present the curves in Figure 5, where $\epsilon_{\mathrm{a}}$ denotes the intensity of unseen attacks and $\epsilon_{\mathrm{i}}$ denotes the intensity of the initial attack for Perturbation Forgery. We can see that the detector trained with a minor $\epsilon_{\mathrm{i}}$ is able to detect adversarial data generated with a larger $\epsilon_{\mathrm{a}}$. When $\epsilon_{\mathrm{i}}$ is larger, adversarial data generated with a minor $\epsilon_{\mathrm{a}}$ is not easy to detect. This might be because adversarial noises generated with a higher $\epsilon_{\mathrm{a}}$ tend to be more obvious to detect. When $\epsilon_{\mathrm{a}}$ is too small than $\epsilon_{\mathrm{i}}$, the distance between the initial attack noise and the unseen attack noise is too far away, reducing the detection performance.
Figure 7: 2D T-SNE visualizations. (a) CIFAR-10 flattened noises of adversarial data and Perturbation Forgery. (b) CIFAR-10 features extracted by the model trained with Perturbation Forgery.
Appendix C Limitations
The major limitation of our method is that finding the optimal scale parameters for distribution perturbation is challenging. We must rely on continuous experimentation to determine these parameters. Specifically, we initiate parameters of uniform distribution as $\epsilon_{\mu}=1$ and $\epsilon_{\sigma}=0.001$, and adjust them using a small step factor. Finally, we find the experiment results when $\epsilon_{\mu}=3$ and $\epsilon_{\sigma}=0.005$ are close to optimal. However, even if these parameters are not optimal, our method still maintains consistent detection performance across various attacks on multiple general and facial datasets, with a satisfactory inference time cost.
Appendix D More Visualization Results
Visualization on CIFAR-10
We extract features and flattened noises from adversarial data and Perturbation Forgery from CIFAR-10 and visualize them using 2D t-SNE projection in Figure 7.
Visualization Examples
Some examples of Perturbation Forgery at $\epsilon=4/255$ are shown in Figure 6.

### v4

@@
-Detecting Adversarial Data via Perturbation Forgery
+Detecting Adversarial Data Using Perturbation Forgery
 Abstract
@@
 Although previous detection methods achieve high performance in detecting gradient-based adversarial attacks, new attacks based on generative models with imbalanced and anisotropic noise patterns evade detection.
-Even worse, existing techniques either necessitate access to attack data before deploying a defense or incur a significant time cost for inference, rendering them impractical for defending against newly emerging attacks that are unseen by defenders.
-In this paper, we explore the proximity relationship between adversarial noise distributions and demonstrate the existence of an open covering for them.
-By learning to distinguish this open covering from the distribution of natural data, we can develop a detector with strong generalization capabilities against all types of adversarial attacks.
-Based on this insight, we heuristically propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production, to train an adversarial detector capable of detecting unseen gradient-based, generative-based, and physical adversarial attacks, while remaining agnostic to any specific models.
+Even worse, the significant inference time overhead and limited performance against unseen attacks make existing techniques impractical for real-world use.
+In this paper, we explore the proximity relationship among adversarial noise distributions and demonstrate the existence of an open covering for these distributions.
+By training on the open covering of adversarial noise distributions, a detector with strong generalization performance against various types of unseen attacks can be developed.
+Based on this insight, we heuristically propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production, to train an adversarial detector capable of detecting any unseen gradient-based, generative-based, and physical adversarial attacks.
 Comprehensive experiments conducted on multiple general and facial datasets, with a wide spectrum of attacks, validate the strong generalization of our method.
+[^1] Code at https://github.com/cc13qq/PFD
 Table 1: Properties of our method vs Other Adversarial Detection: We summarize four metrics for adversarial detection.
-“Detect Attacks” means what kind of attacks the method is designed to detect.
-“Attack-Agnostic” describes whether the detection can detect unseen attack types.
-“Model-Agnostic” describes whether the method is independent of the protected model.
-Finally, “Time Cost” measures the time cost of a detector during inference.
-Methods Detect Attacks Attack-Agnostic Model-Agnostic Time Cost
-LID (Ma et al. 2018) Gradient $\times$ $\times$ $\ast$
-LiBRe (Deng et al. 2021) Gradient $\times$ $\ast$
-[…]
-SPAD (Wang et al. 2023a) Gradient + GAN $\ast$
-Ours Gradient + GAN + Diffusion $\ast$
-Introduction
-Numerous studies have demonstrated the effectiveness of deep neural networks (DNNs) in various tasks (Shi et al.
-2020).
+“Detect Attacks” specifies the types of attacks each method is designed to detect.
+“Model-Agnostic” indicates whether the detection method operates independently of the protected model.
+“Unseen-Attack Detection” assesses the method’s generalization to detect previously unseen attacks, with the reported value representing the AUROC score for detecting the MIFGSM attack on the ImageNet dataset.
+Lastly, “Time Overhead” quantifies the time required for the detector to process 100 samples from the ImageNet dataset.
+Methods Detect Attacks Model-Agnostic Unseen-Attack Detection Time Overhead
+LID [28] Gradient $\times$ 0.9146 1.80s
+LiBRe [8] Gradient $\times$ 0.8725 2.56s
+SPAD [40] Gradient + GAN 0.9820 4.56s
+EPSAD [49] Gradient 0.9918 396.81s
+Ours Gradient + GAN + Diffusion 0.9931 4.85
+1 Introduction
+Numerous studies have demonstrated the effectiveness of deep neural networks (DNNs) in various tasks [36].
 However, it is also well-known that DNNs are vulnerable to adversarial attacks, which generate adversarial data by adding imperceptible perturbations to natural images, potentially causing the model to make abnormal predictions.
 This vulnerability reduces the reliability of deep learning systems, highlighting the urgent need for defense techniques.
-Some approaches employ adversarial training techniques (Zhang et al.
-2019), which incorporate adversarial data into the training process to inherently bolster the model’s immunity against adversarial noise (adv-noise).
-Other methods utilize adversarial purification (Yoon, Hwang, and Lee 2021) to cleanse the input data of adv-noise before model inference.
-Unfortunately, adversarial training often results in diminished classification accuracy on adversarial data and may sacrifice performance on natural data.
+Some approaches employ adversarial training techniques [48], which incorporate adversarial data into the training process to inherently bolster the model’s immunity against adversarial noise (adv-noise).
+Other methods utilize adversarial purification [47] to cleanse the input data of adv-noise before model inference.
+Unfortunately, adversarial training often results in diminished classification accuracy on adversarial data and may sacrifice performance on natural data [1].
 While purification techniques may seem more reliable, the denoising process can inadvertently smooth the high-frequency texture of clean images, leading to suboptimal accuracy on both natural and adversarial images.
-Another branch of adversarial defense is adversarial detection (Feinman et al.
-2017; Ma et al.
-2018), which aims to filter out adversarial data before they are fed into target systems.
+Another branch of adversarial defense is adversarial detection [12, 28], which aims to filter out adversarial data before they are fed into target systems.
 These methods rely on discerning the discrepancies between adversarial and natural distributions, thereby maintaining the integrity of target systems while achieving high performance in defending against adversarial attacks.
-However, existing adversarial detection approaches primarily train specialized detectors tailored for specific attacks or classifiers (Deng et al.
-2021), which may not generalize well to unseen advanced attacks.
-Some recent approaches (Zhang et al.
-2023; Wang et al.
-2023a) have attempted to enhance the generalization of adversarial detectors by leveraging diffusion models and manually designed pseudo-noise.
-Nevertheless, as shown in Table 1, this improvement often comes at the cost of increased inference time, and even well-trained detectors still struggle with adversarial data produced by generative models such as generative adversarial networks (GANs) and diffusion models.
-Specifically, generative-based adversarial attacks, such as M3D (Zhao et al.
-2023) and Diff-PGD (Xue et al.
-2024), are recently developed techniques designed to generate more natural adversarial data.
-Unlike traditional gradient-based attacks like PGD (Madry et al.
-2017), generative-based attacks tend to create imbalanced and anisotropic perturbations that are intense in high-frequency and salient areas but mild in low-frequency and background areas.
-These perturbations circumvent all current adversarial detection methods, posing a new challenge to the field.
+However, existing adversarial detection approaches primarily train specialized detectors tailored for specific attacks or classifiers [8], which may not generalize well to unseen advanced attacks [39, 2].
+Some recent approaches [49, 40] have attempted to enhance the generalization of adversarial detectors by leveraging diffusion models and manually designed pseudo-noise.
+Nevertheless, as shown in Table 1, this improvement often comes at the expense of increased inference overhead, and even well-trained detectors still struggle with adversarial data produced by generative models such as generative adversarial networks (GANs) and diffusion models.
+Specifically, generative-based adversarial attacks, such as M3D [51] and Diff-PGD [44], are recently developed techniques designed to generate more natural adversarial data.
+Unlike traditional gradient-based attacks like PGD [29], generative-based attacks tend to create imbalanced and anisotropic perturbations that are intense in high-frequency and salient areas but mild in low-frequency and background areas.
+These perturbations make them harder to detect with current methods, posing a new challenge to the field of adversarial detection.
 Therefore, there is an urgent need for a low-time-cost detection method with strong generalization performance against both unseen gradient-based attacks and unseen generative-based attacks.
 In this paper, we investigate the association between the distributions of different adv-noises.
-As shown in Figure 1, by modeling noise as multivariant Gaussian distributions, we define the proximity relationship of the noise distributions in a metric space and establish a perturbation method to construct the proximal distributions.
-Based on these assumptions and definitions, we demonstrate that all adv-noise distributions are proximal using the Wasserstein distance, and we deduce a corollary that an open covering of the adv-noise distributions exists.
-We then propose the core idea of this paper: by learning to distinguish the open covering of adv-noise distributions from the natural distribution, we obtain a detector with strong generalization performance against all types of unseen attacks.
-Figure 1: Illustration of the proximity of adv-noise distributions and perturbed noise distributions.
-Left: all the distributions of adv-noise and the perturbed distributions are in a $\epsilon$-ball centered on a given adv-noise distribution.
-Right: by continuously perturbing the commonly used adv-noise distribution, we obtain an open covering of distributions of adversarial noise.
-Based on this idea, we propose Perturbation Forgery, a pseudo-adversarial data generation method designed to train a detector with strong generalization capabilities for detecting all types of adversarial attacks.
+As shown in Figure 1, by modeling noise as multivariate Gaussian distributions, we define the proximity relationship of the noise distributions in a metric space and establish a perturbation method to construct the proximal distributions.
+Based on the assumption and definition, we demonstrate that all adv-noise distributions are proximal using the Wasserstein distance, and we deduce a corollary that an open covering of the adv-noise distributions exists.
+We then propose the core idea of this paper: by training on the open covering, we obtain a detector with strong generalization performance against various types of unseen attacks.
+Figure 1: Illustration of the proximity of adversarial noise distributions and perturbed noise distributions.
+Left: all the distributions of adversarial noise and the perturbed distributions are in a $\varepsilon$-ball centered on a given adversarial noise distribution.
+Right: by continuously perturbing the given adversarial noise distribution, we obtain an open covering of distributions of adversarial noise.
+Based on this idea, we propose Perturbation Forgery, a pseudo-adversarial data (pseudo-adv-data) generation method designed to train a detector with strong generalization capabilities for detecting various types of adversarial attacks.
 As shown in Figure 2, Perturbation Forgery comprises noise distribution perturbation, sparse mask generation, and pseudo-adversarial data (pseudo-adv-data) production.
-By continuously perturbing the noise distribution of a commonly used attack, such as FGSM (Goodfellow, Shlens, and Szegedy 2014), we obtain a nearly complete family of distributions that form the aforementioned open covering.
+By continuously perturbing the noise distribution of a commonly used attack, such as FGSM [15], we obtain a nearly complete family of distributions that form the aforementioned open covering.
 Simultaneously, sparse masks are generated using attention maps and saliency maps of the natural data, converting the sampled global noises into local forms to help the detector focus on local noise patterns so as to detect generative-based adv-noise.
-Sampling noises from the perturbed noise distributions, we sparsify these noises with the sparse masks and inject them into natural data.
-This process converts half of the natural data into pseudo-adv-data, which are then combined with the remaining uncontaminated data to train the binary adversarial detector.
-To the best of our knowledge, we are the first to empower a detector with the robust ability to detect all types of adversarial attacks, including gradient-based, GAN-based, diffusion-based, and physical attacks, with minor time cost.
+By sampling noise from the perturbed noise distributions, we sparsify it using sparse masks and inject it into natural data.
+This process converts half of the natural data into pseudo-adv-data, replacing actual adversarial data.
+The pseudo-adv-data is then combined with the remaining uncontaminated data to train the binary adversarial detector.
+To the best of our knowledge, we are the first to empower a detector with the robust ability to detect various types of adversarial attacks, including gradient-based, GAN-based, diffusion-based, and physical attacks, with minor inference time overhead.
 Our contributions are summarized in four thrusts:
 $\bullet$ By modeling adversarial noise as Gaussian distributions, we investigate the association between different adversarial noise distributions and demonstrate their proximity in a metric space.
-$\bullet$ Based on our assumptions and definitions, we deduce the existence of an open covering for adv-noise distributions and propose training a detector to distinguish this open covering from the natural data distribution, thereby achieving strong generalization.
+$\bullet$ Based on our assumption and definition, we deduce the existence of an open covering for adversarial noise distributions and propose training a detector to distinguish this open covering from the natural data distribution, thereby achieving strong generalization in detection.
 $\bullet$ Building on our deductions, we propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production.
-This method is designed to train a binary detector with strong generalization performance against all types of unseen attacks, including gradient-based, GAN-based, diffusion-based, and physical adversarial data, with minor time cost.
+This method is designed to train a binary detector with strong generalization performance against various types of unseen attacks, including gradient-based, GAN-based, diffusion-based, and physical adversarial data, with minor inference time overhead.
 $\bullet$ We conduct evaluations on multiple general and facial datasets across a wide spectrum of gradient-based, generative-based, and physical attacks, demonstrating the consistently strong performance of our method in adversarial detection tasks.
-Related Work
-Adversarial Attack
+2 Related Work
+2.1 Adversarial Attack
 Gradient-based adversarial attack
-Adversarial attacks (Goodfellow, Shlens, and Szegedy 2014; Madry et al.
-2017; Kurakin, Goodfellow, and Bengio 2016) attempt to fool a classifier into changing prediction results by injecting subtle perturbations to the input data while maintaining imperceptibility from human eyes.
-In an effort to improve transferability under the black-box setting, DIM (Wu et al.
-2021), SIM (Lin et al.
-2019), and VIM (Wang and He 2021) are recent attacks and achieve a high attack rate against several defense techniques.
+Adversarial attacks [15] attempt to fool a classifier into changing prediction results by injecting subtle perturbations to the input data while maintaining imperceptibility from human eyes.
+By introducing iterative perturbation, PGD [29] and BIM [24] generate adversarial data with significantly improved attack performance, while DIM [43], NIM [26], and MIM [9] improve the transferability under the black-box setting.
+Besides, adaptive attacks [6, 13, 18] are proposed to exploit the weak points of each defense, achieving a high attack rate against several defense techniques.
 Generative-based Adversarial Attacks
-Recently, another branch of attack uses generative models to make adversarial data with a more natural style and imbalanced noise compared to gradient-based attacks.
-Among them, CDA (Naseer et al.
-2019), TTP (Naseer et al.
-2021), and M3D (Zhao et al.
-2023) utilize GAN to craft adversarial data and achieve higher efficiency when attacking large-scale datasets.
-Diff-Attack (Chen et al.
-2023) and Diff-PGD (Xue et al.
-2024) employ the diffusion model as a strong prior to better ensure the realism of generated data.
-Designed for perturbing facial recognition systems, AdvMakeup (Yin et al.
-2021) and AMT-GAN (Hu et al.
-2022) are presented to generate imperceptible adversarial images of the face.
-In this paper, a wide spectrum of gradient-based generative-based and adversarial attacks on ImageNet and facial datasets is used to for evaluation.
-Adversarial Detection
-As one of the technical solutions to ensure the safety of DNNs, there has been a significant amount of exploration in the field of adversarial detection (Ma et al.
-2018).
-LiBRe (Deng et al.
-2021) leverages Bayesian neural networks to endow pre-trained DNNs with the ability to defend against unseen adversarial attacks, SPAD (Wang et al.
-2023a) leverages a data augmentation method to detect adversarial face images from unseen attack, while EPSAD (Zhang et al.
-2023) incorporates a diffusion model to detect adversarial data with minimal perturbation magnitude, albeit at the cost of inference time.
-While current methods have achieved remarkable improvements in adversarial detection, generative-based adversarial attacks can easily circumvent them, posing a new challenge.
-To address the problems above, our method is proposed to detect all types of unseen attacks, including gradient-based, generative-based, and physical attacks.
-Proximity of Adversarial Noise Distribution
+Recently, another branch of attacks uses generative models to produce adversarial data with a more natural style and imbalanced noise compared to gradient-based attacks.
+Among them, CDA [31], TTP [30], and M3D [51] utilize GAN to craft adversarial data and achieve higher efficiency when attacking large-scale datasets.
+Diff-Attack [4] and Diff-PGD [44] employ the diffusion model as a strong prior to better ensure the realism of generated data.
+Designed for perturbing facial recognition systems, AdvMakeup [46] and AMT-GAN [19] are presented to generate imperceptible adversarial images of the face.
+In this paper, a wide spectrum of gradient-based, generative-based, and physical adversarial attacks on ImageNet and facial datasets is used to evaluate the detection performance of our method.
+2.2 Adversarial Detection
+As one of the technical solutions to ensure the safety of DNNs, there has been a significant amount of exploration in the field of adversarial detection [28].
+LiBRe [8] leverages Bayesian neural networks to endow pre-trained DNNs with the ability to defend against unseen adversarial attacks, SPAD [40] leverages a data augmentation method to detect adversarial face images from unseen attack, while EPSAD [49] incorporates a diffusion model to detect adversarial data with minimal attack intensity, albeit at the cost of inference time.
+While current methods have achieved remarkable improvements in adversarial detection, generative-based adversarial attacks are harder to detect, posing a new challenge.
+To address the problems above, our method is proposed to detect various types of unseen attacks, including gradient-based, generative-based, and physical attacks.
+3 Proximity of Adversarial Noise Distribution
 In this section, we provide an overview of preliminary concepts related to adversarial data generation and explore the relationship between adversarial noise distributions.
-We define the proximity of noise distributions and examine the imperceptibility of adversarial noise.
-Following this, we present the core idea of this paper: by learning to distinguish the open covering of adversarial noise distributions from the natural distribution, we develop a detector with strong generalization capabilities against all types of unseen attacks.
-Adversarial data generation
-Given a DNN $f$ that works on dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}^{n}_{i=1}$ with $(\mathbf{x}_{i}\in\mathcal{X},y_{i}\in\mathcal{C})$ being (sample, ground-truth) pair, an adversarial example $\hat{\mathbf{x}}$ regarding sample $\mathbf{x}$ with perturbation magnitude $\epsilon$ is calculated as:
-$\hat{\mathbf{x}}=\mathop{\mathrm{argmax}}_{\hat{\mathbf{x}}\in\mathbb{B}(\mathbf{x},\epsilon)}\mathcal{L}(f(\hat{\mathbf{x}}),y),$ (1)
+We define the proximity of noise distributions and present the core idea of this paper: by training on the open covering of adversarial noise distributions, we obtain a detector with strong generalization performance against various types of unseen attacks.
+3.1 Adversarial data generation
+Given a DNN $f$ that works on dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}^{n}_{i=1}$ with $(\mathbf{x}_{i}\in\mathcal{X},y_{i}\in\mathcal{C})$ being (sample, ground-truth) pair, an adversarial example $\hat{\mathbf{x}}$ regarding sample $\mathbf{x}$ with attack intensity $\epsilon$ is calculated as:
+$\hat{\mathbf{x}}=\mathop{\mathrm{argmax}}_{\hat{\mathbf{x}}\in\mathcal{B}(\mathbf{x},\epsilon)}\mathcal{L}(f(\hat{\mathbf{x}}),y),$ (1)
 where $\mathcal{L}$ is some loss function and $\mathcal{B}(\mathbf{x},\epsilon)=\{\mathbf{x^{\prime}}|\ell^{p}(\mathbf{x},\mathbf{x^{\prime}})<\epsilon\}$ restricts the perturbation under $\ell^{p}$ norm.
-Following (Goodfellow, Shlens, and Szegedy 2014), we simply denote adversarial sample as $\hat{\mathbf{x}}=\mathbf{x}+\boldsymbol{\eta}$.
-Proximity of noise distribution
+Following [15], we simply denote adversarial sample as $\hat{\mathbf{x}}=\mathbf{x}+\boldsymbol{\eta}$.
+Figure 2: Illustration of Perturbation Forgery.
+Before training, we estimate the noise distribution from a commonly used attack, then continuously perturb it in each batch to create an open covering of the adversarial noise distributions.
+Next, noises sampled from these perturbed distributions are converted into localized noise by applying sparse masks.
+Finally, pseudo-adversarial data is generated by adding these localized noises to natural samples.
+3.2 Proximity of noise distribution
 Due to the subtle and imperceptible nature of adv-noise, adversarial data and natural data appear visually similar.
 Thus, we assume that this similarity also exists between the noises of different adversarial attacks.
-To investigate the similarity of these noises, we provide the following definition, which depicts a close relationship between two independent noise distributions:
+To investigate the similarity of these noises, we provide the following definition, which depicts a close relationship between two noise distributions.
 Definition 1 (Proximal Noise Distribution).
@@
 We call them proximal noise distributions under $\varepsilon$ if we have a metric $d(\cdot,\cdot)$ and $\exists\ \varepsilon>0$, s.t.
-$d(P,Q)<\varepsilon.$ (2)
-A common assumption for instance feature distribution is to model it as a Gaussian distribution (Du et al.
-2022).
-Obtained by subtracting from clean data and adversarial data, adv-noises naturally form a Gaussian distribution.
-Therefore, we assume that noise r.v $\boldsymbol{\eta}$ satisfied truncated Gaussian distribution $\mathcal{N}(\mu,\sigma^{2}\mathbf{I}_{d})$ with it’s truncated interval is $[-\epsilon,\epsilon]^{d}$.
+$d(P,Q)<\varepsilon$.
+A common assumption for instance feature distribution is to model it as a Gaussian distribution [10].
+Obtained by subtracting from natural data and adversarial data, adv-noises naturally form a Gaussian distribution.
+Therefore, we assume that noise r.v $\boldsymbol{\eta}\in\mathbb{R}^{d}$ satisfied truncated Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\sigma^{2}\mathbf{I}_{d})$ with it’s truncated interval is $[-\epsilon,\epsilon]^{d}$.
 Built upon Definition 1, we derive the following theorem about the association between the distributions of different adv-noises:
@@
 For $\forall\ i\neq j$, $Q_{i}$ and $Q_{j}$ are proximal noise distributions if the following conditions are met.
-1) if $Q_{i},Q_{j}\in\mathcal{P}_{a}$.
+1) $Q_{i},Q_{j}\in\mathcal{P}_{a}$.
 […]
-Theorem 1 tells us: 1) All adversarial noise distributions are proximal noise distributions.
+Theorem 1 tells us: 1) All the adversarial noise distributions under the same $\ell^{p}$ norm are proximal noise distributions.
 2) If the parameters of two noise distributions are sufficiently close to those of a given distribution, we can regard these two distributions as proximal noise distributions.
-Figure 2: Illustration of Perturbation Forgery.
-Before training, we estimate the noise distribution of a commonly used attack and continuously perturb it in each batch.
-After that, we sample noises from the perturbed distribution and convert them into local noise using sparse masks.
-Finally, pseudo-adversarial data is generated by adding these local noises to the natural samples.
 We can further deduce several corollaries of the theorem.
@@
 2) The adv-noise distribution set is also located in this ball: $\mathcal{P}_{a}\subset\mathcal{P}_{\varepsilon}$.
-3) Based on the properties of the metric space $(\mathcal{P}_{\varepsilon},d)$, for all subset $\mathcal{P}_{i}\subset\mathcal{P}_{\varepsilon}$, there exists an open covering of the subset: $\mathcal{P}_{i}\subset\{G_{c}\}_{c\in J}$.
-Based on the corollaries, we deduce the core idea of this paper: Let $p(\hat{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})=\mathcal{N}(\mu+\epsilon_{\mu i},(\sigma+\epsilon_{\sigma i})^{2}\mathbf{I}_{d}),i=1,2,\dots,n$ denote generated proximal noise distributions of the given adv-noise distribution $p(\boldsymbol{\eta})=\mathcal{N}(\mu,\sigma^{2}\mathbf{I}_{d})$ where $\epsilon_{\mu i},\epsilon_{\sigma i}$ are randomly sampled from uniform distributions and restricted by $\epsilon_{\mu i},\epsilon_{\sigma i}<\varepsilon$.
+3) Based on the properties of the metric space $(\mathcal{P}_{\varepsilon},d)$, for all subset $\mathcal{P}_{i}\subset\mathcal{P}_{\varepsilon}$, there exists an open covering of the subset: $\mathcal{P}_{i}\subset\{G_{c}\}_{c\in\mathbb{I}}$.
+Based on the corollaries, we deduce the core idea of this paper: Let $p(\hat{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})=\mathcal{N}(\boldsymbol{\mu}+\epsilon_{\mu i},(\sigma+\epsilon_{\sigma i})^{2}\mathbf{I}_{d}),i=1,2,\dots,n$ denote generated proximal noise distributions of the given adv-noise distribution $p(\boldsymbol{\eta})=\mathcal{N}(\boldsymbol{\mu},\sigma^{2}\mathbf{I}_{d})$, where $\epsilon_{\mu i},\epsilon_{\sigma i}$ are randomly sampled from uniform distributions and restricted by $\epsilon_{\mu i},\epsilon_{\sigma i}<\varepsilon$.
 As shown in Figure 1, when $n\rightarrow\infty$, we can always obtain a distribution set $\mathcal{P}_{c}=\{p(\hat{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})\}_{i=1}^{\infty}$ that forms an open covering of the adversarial distribution set $\mathcal{P}_{a}$.
-Trained with the open covering, the detector can distinguish the distributions of natural data and adversarial data, thereby having the functionality to detect unseen adversarial attacks.
-For proofs of theorem and corollaries, please refer to the supplementary materials.
-Perturbation Forgery
-Based on the theory support in the last section, we heuristically propose Perturbation Forgery to train a detector with strong generalization against all types of unseen attacks.
-As illustrated in Figure 2, Perturbation Forgery encompasses three elements: distribution perturbation, sparse mask generation, and pseudo-adversarial data production.
+By training on the open covering, the detector learns to distinguish between the distributions of natural and adversarial data, enabling it to effectively detect unseen adversarial attacks.
+Analysis of distribution assumption, proofs of theorem, and corollaries are provided in the supplementary materials.
+4 Perturbation Forgery
+Based on the theory in the last section, we heuristically propose Perturbation Forgery to train a detector with strong generalization against various types of unseen attacks.
+As illustrated in Figure 2, Perturbation Forgery encompasses three elements: noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production.
 At first, the noise distribution of a commonly used attack is continuously perturbed using noise distribution perturbation into numerous noise distributions that remain proximal to the original, thereby forming the open covering.
 After that, we sample noises from the perturbed distribution and convert them into sparse local noise using sparse mask generation.
-Subsequently, we convert half of the natural data into pseudo-adversarial data using Pseudo-Adversarial Data Generation and train a binary detector on the these data along with the remaining natural data.
-Noise Distribution Perturbation
+Subsequently, we convert half of the natural data into pseudo-adv-data using pseudo-adversarial data production.
+This pseudo-adv data replaces actual adversarial data and is combined with the remaining natural data to train a binary detector.
+4.1 Noise Distribution Perturbation
 Given the training dataset $\{(\mathbf{x}_{k},y_{k})\}^{n}_{k=1}$ composed by natural data, we add adv-noise on it using a commonly used attack (such as FGSM) and generate adversarial noise features $\mathcal{D}_{\mathrm{noise}}:=\{\mathbf{z}_{k}=h(\boldsymbol{\eta}_{k})\}_{k=1}^{n}$ by
@@
 where $h:\mathcal{X}\rightarrow\mathbb{R}^{d}$ is the flattening operation that converts multidimensional data into vectors.
-Assuming the noise features form a multivariate Gaussian distribution $p(\mathbf{z})=\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ (Du et al.
-2022), we estimate the parameters mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$ of $\mathcal{D}_{\mathrm{noise}}$ by
+Assuming the noise features form a multivariate Gaussian distribution $p(\mathbf{z})=\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ [10], we estimate the parameters mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$ of $\mathcal{D}_{\mathrm{noise}}$ by
 $\widehat{\boldsymbol{\mu}}=\frac{1}{N}\sum_{k}\mathbf{z}_{k},\quad\widehat{\boldsymbol{\Sigma}}=\frac{1}{N}\sum_{k}\left(\mathbf{z}_{k}-\widehat{\boldsymbol{\mu}}\right)\left(\mathbf{z}_{k}-\widehat{\boldsymbol{\mu}}\right)^{\top}.$ (3)
@@
 where $\boldsymbol{m}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ and $\boldsymbol{V}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d\times d})$ are vector and matrix of perturbation.
-$\alpha_{i}\sim\boldsymbol{U}(-\epsilon_{\mu},\epsilon_{\mu})$ and $\beta_{i}\sim\boldsymbol{U}(0,\epsilon_{\sigma})$ are scale parameters sampled from uniform distributions.
-After $n$ batches of training, we have a perturbed distribution set $\{\mathcal{N}(\widehat{\boldsymbol{\mu}}_{i},\widehat{\boldsymbol{\Sigma}}_{i})\}$, thereby forming the approximate open covering $\mathcal{P}_{c}$.
+$\alpha_{i}\sim\boldsymbol{U}(-\epsilon_{\mu},\epsilon_{\mu})$ and $\beta_{i}\sim\boldsymbol{U}(0,\epsilon_{\sigma})$ are scale parameters.After $n$ batches of training, we have a perturbed distribution set $\{\mathcal{N}(\widehat{\boldsymbol{\mu}}_{i},\widehat{\boldsymbol{\Sigma}}_{i})\}$, thereby forming the approximate open covering $\mathcal{P}_{c}$.
 Trained with $\mathcal{P}_{c}$, the detector can distinguish the distributions of natural and adversarial data.
-Sparse Mask Generation
-Unlike gradient-based adversarial attacks, generative-based attacks tend to generate imbalanced and anisotropy perturbations, showing much intensity in high-frequency and salient areas and mild in low-frequency and background areas.
-Besides, in the physical situation, adversarial attacks also tend to add patches to local areas.
+4.2 Sparse Mask Generation
+Unlike gradient-based adversarial attacks, generative-based attacks tend to generate imbalanced and anisotropic perturbations, showing much intensity in high-frequency and salient areas and mild in low-frequency and background areas.
+Besides, in physical scenarios, adversarial attacks also tend to add patches to local areas.
 Noticing this, we propose sparse masks to perturbation forgery to assist the detector in detecting local adv-noises.
-Previous research has demonstrated that the spectra of real and fake images differ significantly in high-frequency regions (Luo et al.
-2021).
-Additionally, generative-based adversarial data manipulate characteristics, such as abnormal color aberrations, in these high-frequency regions (Wang et al.
-2023a).
+Previous research has demonstrated that the spectra of real and fake images differ significantly in high-frequency regions [27].
+Additionally, generative-based adversarial data manipulate characteristics, such as abnormal color aberrations, in these high-frequency regions [40].
 To augment the detection performance on generative-based adversarial data, we need to sparse the pseudo-noise sampled from perturbed noise distribution and highlight the noise in high-frequency regions to mimic the style of generative-based attacks.
 Algorithm 1 Training Procedure with Perturbation Forgery
-Input: Training dataset, initial attack $\operatorname{Attack}$, saliency detection model $f_{s}$, and GradCAM model $f_{c}$.
-Output: A detector with strong generalization capabilities against all types of unseen attacks.
-1:  Generate adversarial noise using the initial attack $\operatorname{Attack}$ in Eq. 3.
+Input: Training dataset, a commonly used attack, saliency detection model $f_{s}$, and GradCAM model $f_{c}$.
+Output: A detector with strong generalization capabilities against various types of unseen attacks.
+1:  Generate adversarial noise using the commonly used attack in Eq. 2.
 2:  Estimate a multivariate Gaussian distribution $p(\mathbf{z})=\mathcal{N}(\widehat{\boldsymbol{\mu}},\widehat{\boldsymbol{\Sigma}})$ of adv-noise using Eq. 3.
@@
 4:     Generate a perturbed noise distribution $p(\mathbf{z}_{i})=\mathcal{N}(\widehat{\boldsymbol{\mu}}_{i},\widehat{\boldsymbol{\Sigma}}_{i})$ using Eq. 4.
-5:     for sample $\mathbf{x}_{i,k}$ half of batch $i$ do
+5:     for sample $\mathbf{x}_{i,k}$ in half of batch $i$ do
 6:        Generate sparse masks $\operatorname{Mask}(\mathbf{x}_{i,k})$ using Eq. 8.
-7:        Sample pseudo-adversarial noise $\widehat{\boldsymbol{\eta}}_{i,k}$ using Eq. 10.
-8:        Generate pseudo-adversarial data $\widehat{\mathbf{x}}_{i,k}$ using Eq. 12.
+7:        Sample pseudo-adv noise $\widehat{\boldsymbol{\eta}}_{i,k}$ using Eq. 9.
+8:        Generate pseudo-adv data $\widehat{\mathbf{x}}_{i,k}$ using Eq. 11 and replace $\mathbf{x}_{i,k}$.
 9:     end for
@@
 11:  end for
-Given this, we first mask the high-frequency area of a natural sample using saliency detection (Qin et al. 2019) and GradCAM (Selvaraju et al. 2017):
-$\operatorname{Mask}_{1}(\mathbf{x})=\operatorname{Map}(f_{s}(\mathbf{x});\gamma_{s})\cup\operatorname{Map}(f_{c}(\mathbf{x});\gamma_{c}),$ (6)
+Given this, for a natural sample $\mathbf{x}_{i,k}$ in the $\text{i}^{th}$ batch, we first mask the high-frequency area using saliency detection [33] and GradCAM [34]:
+$\operatorname{Mask}_{1}(\mathbf{x}_{i,k})=\operatorname{Map}(f_{s}(\mathbf{x}_{i,k});\gamma_{s})\cup\operatorname{Map}(f_{c}(\mathbf{x}_{i,k});\gamma_{c}),$ (5)
 where $f_{s}$ and $f_{c}$ are the saliency detection model and GradCAM model respectively. $\operatorname{Map}(\mathbf{x};\gamma)$ is a binary indicator function parameterized by $\gamma$:
-$\operatorname{Map}(\mathbf{x};\gamma)=\operatorname{Clip}_{[0,1]}(\mathbf{x}-\gamma\mathbf{I}_{\mathcal{X}}).$ (7)
-The noise that occurs in the high-frequency region tends to occupy only a small part of the region rather than spreading to the entire high-frequency spectrum.
-Therefore, we select the high-frequency point using the Sobel operator (Duda and Hart 1973) to sparsify the mask of sample $\mathbf{x}_{i,k}$:
-$\operatorname{Mask}_{2}(\mathbf{x})=\operatorname{Map}(\operatorname{Sobel}(\mathbf{x});\gamma_{l}),$ (8)
+$\operatorname{Map}(\mathbf{x}_{i,k};\gamma)=\operatorname{Clip}_{[0,1]}(\mathbf{x}_{i,k}-\gamma\mathbf{I}_{\mathcal{X}}).$ (6)
+The adv-noise that occurs in the high-frequency region tends to occupy only a small part of the region.
+Therefore, we select the high-frequency point from the gradient map obtained by Sobel operator [11] to sparsify the mask of sample $\mathbf{x}_{i,k}$:
+$\operatorname{Mask}_{2}(\mathbf{x}_{i,k})=\operatorname{Map}(\operatorname{Sobel}(\mathbf{x}_{i,k});\gamma_{l}),$ (7)
 $\operatorname{Mask}(\mathbf{x}_{i,k})=\operatorname{Mask}_{1}(\mathbf{x}_{i,k})\cap\operatorname{Mask}_{2}(\mathbf{x}_{i,k}).$ (8)
-Pseudo-Adversarial Data Production
-For the $i^{\mathrm{th}}$ batch of training, we sample pseudo-adversarial noise from the perturbed distribution by:
+4.3 Pseudo-Adversarial Data Production
+For the $i^{\mathrm{th}}$ batch of training, we sample pseudo-adv noise from the perturbed distribution by:
 $\widehat{\boldsymbol{\eta}}_{i,k}=h^{-1}(\mathbf{z}_{i,k}),$ (9)
@@
 Note that $\boldsymbol{\eta}_{i}$ is a global noise and needs to be localized by a sparse mask.
-For natural data $\mathbf{x}_{i,k}$ in the $i^{\mathrm{th}}$ batch, the corresponding pseudo-adversarial data $\widehat{\boldsymbol{x}}_{i,k}$ is produced by adding the masked local noise to it:
+The pseudo-adv data $\widehat{\mathbf{x}}_{i,k}$ is produced by adding the masked local noise to $\mathbf{x}_{i,k}$:
 $\widehat{\mathbf{x}}_{i,k}=\mathbf{x}_{i,k}+\widehat{\boldsymbol{\eta}}_{i,k}\otimes\operatorname{Mask}(\mathbf{\mathbf{x}_{i,k}}).$ (11)
 where $\otimes$ is Hadamard product.
-For each batch, we convert half of the natural data to pseudo-adversarial data by Perturbation Forgery and train a binary detector using the Cross-Entropy classification loss (Zhang and Sabuncu 2018).
+For each batch, we convert half of the natural data to pseudo-adv data by Perturbation Forgery and train a binary detector using the Cross-Entropy classification loss [50].
 The overall training procedure with Perturbation Forgery is presented in Algorithm 1.
-Experiments
+Table 2: Comparison of AUROC scores of detecting gradient-based and adaptive attacks on ImageNet100 under $\epsilon=4/255$.
+Detector BIM PGD DIM MIM NIM AA BPDA+EOT MM
+LID [28] 0.9782 0.9750 0.8942 0.9146 0.8977 0.9124 0.9202 0.9227
+LiBRe [8] 0.9259 0.9548 0.9243 0.8725 0.9013 0.8653 0.8714 0.8573
+SPAD [40] 0.9846 0.9851 0.9815 0.9820 0.9823 0.9890 0.9885 0.9811
+EPSAD [49] 0.9998 0.9989 0.9923 0.9918 0.9972 0.9998 0.9985 0.9923
+ours 0.9911 0.9912 0.9863 0.9931 0.9934 0.9941 0.9927 0.9944
+5 Experiments
 To validate the detection performance of our approach against various adversarial attack methods, we conduct extensive experiments on multiple general and facial datasets.
-Experiment Settings
-Datasets and Network Architectures
-The datasets for evaluation include 1) general dataset ImageNet100 (Deng et al.
-2009), and 2) face datasets: Makeup (Gu et al.
-2019), CelebA-HQ (Karras et al.
-2017), and LFW (Gary et al.
-2007).
-Gradient-based attacks and generative-based attacks are launched only on ImageNet100.
-For face attacks, we conduct Adv-Makeup on Makeup, AMT-GAN on CelebA-HQ, and physical attacks on LFW.
+For additional experiments, analyses, and a discussion of limitations, please refer to the supplementary material.
+5.1 Experiment Settings
+Datasets
+The datasets used for evaluation include: 1) general datasets: ImageNet100 [7] and CIFAR-10 [23] for evaluate gradient-based and generative-based attacks, and 2) face datasets: Makeup [16] for Adv-Makeup [46], CelebA-HQ [21] for AMT-GAN [19], and LFW [14] for TIPIM [45] and physical attacks.
+Table 3: Comparison of AUROC scores of detecting generative-based attacks on ImageNet100.
+      Detector       CDA       TTP       M3D       Diff-Attack       Diff-PGD
+      CNN-Detection [41]       0.7051       0.6743       0.6917       0.3963       0.5260
+      LGrad [37]       0.6144       0.6077       0.6068       0.5586       0.5835
+      Universal-Detector [32]       0.7945       0.8170       0.8312       0.9202       0.5531
+      DIRE [42]       0.8976       0.9097       0.9129       0.9097       0.9143
+      SPAD       0.9385       0.9064       0.8984       0.8862       0.8879
+      EPSAD       0.9674       0.6997       0.7265       0.4700       0.1507
+      ours       0.9878       0.9678       0.9364       0.9223       0.9223
+Table 4: Comparison of AUROC scores for detecting various adversarial attacks on the face dataset.
+Detector TIPIM Adv-Sticker Adv-Glasses Adv-Mask Detector Adv-Makeup AMT-GAN
+LID 0.6974 0.5103 0.6472 0.5319 He et al. [17] 0.5252 0.8823
+LiBRe 0.5345 0.9739 0.8140 0.9993 Luo et al. [27] 0.6178 0.6511
+EPSAD 0.9647 0.9462 0.9260 0.9578 ODIN[25] 0.6325 0.6970
+SPAD 0.9256 0.9975 0.9244 0.9863 SPAD 0.9657 0.8969
+ours 0.9885 0.9999 0.9999 0.9999 ours 0.9762 0.9216
 Adversarial Attacks
-We adopt 9 gradient-based attacks, 5 generative-based attacks, and 6 face attacks to evaluate our method.
-For gradient-based attacks, BIM (Kurakin, Goodfellow, and Bengio 2016), PGD (Madry et al.
-2017), RFGSM (Tramèr et al.
-2018), MIM (Dong et al.
-2018), DIM (Wu et al.
-2021), NIM and SNIM (Lin et al.
-2019), VNIM (Wang and He 2021), and Auto-Attack (Croce and Hein 2020) are used as the unseen attacks.
-For generative-based attacks, we adopt 3 GAN-based attacks including CDA (Naseer et al.
-2019), TTP (Naseer et al.
-2021), and M3D (Zhao et al.
-2023), and 2 Diffusion-based attacks including Diff-Attack (Chen et al.
-2023) and Diff-PGD (Xue et al.
-2024) to craft adv-samples.
-For face attacks, 1 gradient-based attack TIPIM (Yang et al.
-2021), 2 GAN-based attacks including AdvMakeup (Yin et al.
-2021) and AMT-GAN (Hu et al.
-2022), and 3 physical attacks including Adv-Sticker (Tong et al.
-2021), Adv-Glasses (Sharif et al.
-2016), and Adv-Mask (Tong et al.
-2021) produce adversarial face samples.
-For each of the above attack methods, adversarial data is generated with the perturbation magnitude of $\epsilon=4/255$ (except for the unrestricted attacks) following (Zhang et al.
-2023).
-Table 2: Comparison of AUC scores (%) of detecting gradient-based attacks on ImageNet100.
-Detector BIM PGD RFGSM DIM MIM NIM VNIM SNIM AA
-LID (Ma et al. 2018) 0.9782 0.9750 0.9637 0.8942 0.9146 0.8977 0.8865 0.8752 0.9124
-LiBRe (Deng et al. 2021) 0.9259 0.9548 0.9769 0.9243 0.8725 0.9013 0.8862 0.8694 0.8653
-SPAD (Wang et al. 2023a) 0.9846 0.9851 0.9833 0.9815 0.9820 0.9823 0.9891 0.9877 0.9890
-EPSAD (Zhang et al. 2023) 0.9998 0.9989 0.9992 0.9923 0.9918 0.9972 0.9946 0.9908 0.9998
-ours 0.9911 0.9912 0.9911 0.9863 0.9931 0.9934 0.9878 0.9914 0.9941
-Table 3: Comparison of AUC scores of detecting generative-based attacks on ImageNet100.
-Detector CDA TTP M3D Diff-Attack Diff-PGD
-CNN-Detection (Wang et al. 2020) 0.7051 0.6743 0.6917 0.3963 0.526
-LGrad (Tan et al. 2023) 0.6144 0.6077 0.6068 0.5586 0.5835
-Universal-Detector (Ojha, Li, and Lee 2023) 0.7945 0.8170 0.8312 0.9202 0.5531
-DIRE (Wang et al. 2023b) 0.8976 0.9097 0.9129 0.9097 0.9143
-EPSAD 0.9674 0.6997 0.7265 0.4700 0.1507
-SPAD 0.9385 0.9064 0.8984 0.8862 0.8879
-ours 0.9878 0.9678 0.9364 0.9223 0.9223
+We adopt 5 gradient-based attacks, 3 adaptive attacks, 5 generative-based attacks, and 6 face attacks to evaluate our method.
+For gradient-based attacks, BIM [24], PGD [29], MIM [9], DIM [43], and NIM [26] are used as the unseen attacks.
+For adaptive attacks, we adopt AutoAttack (AA) [6], BPDA+EOT [18], and Minimum-Margin Attack (MM) [13].
+For generative-based attacks, we adopt 3 GAN-based attacks including CDA [31], TTP [30], and M3D [51], and 2 Diffusion-based attacks including Diff-Attack [4] and Diff-PGD [44].
+For face attacks, we adopt 1 gradient-based attack TIPIM [45], 2 GAN-based attacks including AdvMakeup [46] and AMT-GAN [19], and 3 physical attacks including Adv-Sticker [38], Adv-Glasses [35], and Adv-Mask [38].For each of the above attack methods, adversarial data is generated with the attack intensity of $\epsilon=4/255$ (except for the unrestricted attacks), and iterative attacks run for 10 steps using step size 2.
 Baselines
-We compare our method with state-of-the-art detection methods across three research fields, including adversarial detection, synthetic image detection, and face forgery detection, evaluating the detection of various types of adversarial attacks.
-For detecting gradient-based attacks, we compared our method with adversarial detection methods including LID (Ma et al.
-2018), LiBRe (Deng et al.
-2021), SPAD (Wang et al.
-2023a), and EPSAD (Zhang et al.
-2023).
-Since existing adversarial detection methods do not address generative-based adversarial data, we compared our method with synthetic image detection methods including CNN-Detection(Wang et al.
-2020), LGrad(Tan et al.
-2023), Universal-Detector(Ojha, Li, and Lee 2023), and DIRE(Wang et al.
-2023b) for detecting generative-based attacks.
-For detecting GAN-based face attacks, we compared with face forgery detection methods including Luo et al.
-(Luo et al.
-2021), He et al.
-(He et al.
-2021), and ODIN (Liang, Li, and Srikant 2017).
-We conducted experiments using the pre-trained models and authors’ implementation code of the above methods.
+To evaluate our method across various types of adversarial attacks, we compare it with state-of-the-art detection methods from three related research domains: adversarial detection, synthetic image detection, and face forgery detection.
+For gradient-based attacks, we benchmark against four adversarial detection methods: LID [28], LiBRe [8], SPAD [40], and EPSAD [49].
+Since most adversarial detection methods are not designed to address generative-based adversarial attacks, we include a comparison with four synthetic image detection techniques: CNN-Detection [41], LGrad [37], Universal-Detector [32], and DIRE [42].
+For detecting GAN-based face attacks, we also compare against two face forgery detection methods, Luo et al.[27] and He et al.[17], as well as ODIN [25].
 Evaluation Metrics
-Following previous adversarial detection works (Zhang et al.
-2023), we employ the Area Under the Receiver Operating Characteristic Curve (AUC) as the primary evaluation metric to assess the performance of our classification models.
-The AUC is a widely recognized metric in the field of machine learning and is particularly useful for evaluating binary classification problems.
+Following previous adversarial detection works [49], we employ the Area Under the Receiver Operating Characteristic Curve (AUROC) [20] as the primary evaluation metric to assess the performance of our classification models.
+The AUROC is a widely recognized metric in the field of machine learning and is particularly useful for evaluating binary classification problems.
 For fair comparisons, the number of natural data and adversarial data in the experiments remained consistent.
 Implementation details
-XceptionNet (Chollet 2017) is selected as our backbone model and is trained on all the above datasets.
-For adv-noise distribution estimation, we utilize Torchattacks (Kim 2020) to generate 1000 noise images for each dataset.
+XceptionNet [5] is selected as our backbone model and is trained on all the above datasets.
+For adversarial noise distribution estimation, we utilize Torchattacks [22] to generate 1000 noise images for each dataset.
 The uniform distribution parameters are set as $\epsilon_{\mu}=3$ and $\epsilon_{\sigma}=0.005$.
@@
 The thresholds of the binary indicator function are set as $\gamma_{s}=0$, $\gamma_{c}=125$ and $\gamma_{l}=100$.
-FGSM (Goodfellow, Shlens, and Szegedy 2014) is used as the initial attack for Perturbation Forgery.
+FGSM [15] is used as the initial attack, i.e.
+the commonly used attack, for Perturbation Forgery.
 Training epochs are set to 10 and convergence is witnessed.
-Detecting Gradient-based Attacks
+5.2 Detecting Gradient-based Attacks
 We begin by comparing our method with state-of-the-art techniques in detecting gradient-based attacks on ImageNet100.
-All the attacks are seen to LID but unseen to other baselines and our method.
-For each attack, 1,000 adversarial images are generated in the white-box setting to test the detection models.
+All the attacks are unseen by both the baseline methods and our method.
+For each attack, 1,000 adversarial images are generated in the gray-box setting, where the attacker has access to the victim model but is blind to the detection model.
 All baseline methods are sourced from existing adversarial detection research.
-As reported in Table 2, although EPSAD performs well, our method demonstrates effective detection (achieving an AUC score of 0.9931 against MIM and 0.9914 against SNIM) across a variety of attack methods.
+As reported in Table 2, although EPSAD performs well, our method demonstrates effective detection (achieving an AUROC score of 0.9931 against MIM and 0.9944 against MM) across a variety of attack methods.
 These results validate our assumptions about the proximity of the noise distribution.
@@
 The consistent detection performance verifies the effectiveness of our core idea of training a detector to distinguish the open covering of adv-noise distributions from natural ones and demonstrates the strong generalization capability of our method against gradient-based attacks.
-Additional results on CIFAR10 are provided in the supplementary materials.
-Table 4: Comparison of AUC scores for detecting various adversarial attacks on the face.
-Detector TIPIM Adv-Sticker Adv-Glasses Adv-Mask Detector Adv-Makeup AMT-GAN
-LID 0.6974 0.5103 0.6472 0.5319 He et al. (He et al. 2021) 0.5252 0.8823
-LiBRe 0.5345 0.9739 0.8140 0.9993 Luo et al. (Luo et al. 2021) 0.6178 0.6511
-EPSAD 0.9647 0.9462 0.9260 0.9578 ODIN(Liang, Li, and Srikant 2017) 0.6325 0.6970
-SPAD 0.9256 0.9975 0.9244 0.9863 SPAD 0.9657 0.8969
-ours 0.9885 0.9999 0.9999 0.9999 ours 0.9762 0.9216
-Detecting generative-based Attacks
-To investigate the advantage of our method in detecting generative-based attacks, we compare with state-of-the-art methods of synthetic image detection and adversarial detection methods on ImageNet100.
-As shown in Table 3, our method substantially surpasses other methods, exhibiting highly effective detection capabilities.
-For adversarial detection baselines, due to the noise produced by generative models tending to be imbalanced compared to perturbations of traditional gradient-based attacks, these methods struggle to conduct effective detection.
-In contrast, with the assistance of sparse masks, our detector can perceive the local noises brought by GAN and diffusion models.
-As adversarial data is crafted from natural data, most synthetic image detection methods lose their effectiveness except DIRE(Wang et al.
-2023b).
-This is probably because DIRE leverages the reconstruction error of images generated by diffusion models, thus maintaining a certain level of effectiveness.
-Table 5: Ablation study on masks.
+5.3 Detecting Generative-based Attacks
+To investigate our method’s advantage in detecting generative-based attacks, we compared it with state-of-the-art synthetic image detection and adversarial detection methods on ImageNet100.
+As shown in Table 3, our method significantly outperforms the baseline methods, demonstrating highly effective detection capabilities.
+For adversarial detection baselines, the imbalanced noise generated by GAN and diffusion models makes it difficult for these methods, which are typically designed for gradient-based attacks, to detect effectively.
+In contrast, with the help of sparse masks, our detector is able to detect the localized noise introduced by GANs and diffusion models.
+Most synthetic image detection methods are less effective in detecting adversarial data crafted from natural samples, with the exception of DIRE[42].
+This effectiveness may stem from DIRE’s use of reconstruction error from images generated by diffusion models, which maintains its robustness in this setting.
+5.4 Detecting Face Adversarial Examples
+Adversarial attacks targeting face recognition models represent an important branch of adversarial example research.
+Therefore, the detection ability of adversarial face examples is an important indicator of the performance of the detector.
+We perform various adversarial face attacks and compare our approach with state-of-the-art methods from adversarial face detection and face forgery detection.
+As shown in Table 4, our method shows excellent performance against a variety of adversarial attacks targeting faces and surpasses many existing adversarial face detection techniques.
+It further indicates the ability of the sparse mask to capture the essential characteristics of face adversarial examples and the strong generalization capability of our method.
+Table 5: Ablation study on masks (AUROC score), where $f_{s}$ denotes saliency detection model and $f_{c}$ represents the CAM model.
 Ablation CDA TTP M3D Diff-Attack Diff-PGD
@@
 w/o $\operatorname{Sobel}$ 0.9868 0.7235 0.6984 0.5231 0.5547
-Table 6: Time cost of inference (second) on ImageNet100
-Detector LID LiBRe EPSAD SPAD ours
-Time (second) 1.80 2.56 396.81 4.56 4.85
-Detecting Face Adversarial Examples
-Adversarial attack on the face recognition model is an important branch of adversarial example research.
-Therefore, the detection ability of adversarial face examples is an important indicator of the performance of the detector.
-We perform various adversarial face attacks and compare our approach with state-of-the-art methods from adversarial face detection and face forgery detection.
-As shown in Table 4, our method shows excellent performance against a variety of adversarial attacks targeting faces and surpasses many existing adversarial face detection techniques.
-It further indicates the ability of the sparse mask to capture the essential characteristics of face adversarial examples and the strong generalization capability of our method.
-Ablation Study and Analysis
-To further evaluate our method, we conduct experiments to show the effect of the sparse mask and compare the time cost to baseline methods.
-We also analyze the functionality of Perturbation Forgery through visualizations.
-The impact of hyper-parameters is reported in the supplementary materials.
-Impact of each mask
-As argued, the sparse mask is proposed to assist the detector in identifying local adversarial noises present in generative-based adversarial data.
+5.5 Ablation Study and Analysis
+To further evaluate our method, we report an ablation study on the sparse mask, white-box attack detection results, the impact of the initial attack and the data size for distribution estimation, and the functionality of Perturbation Forgery through visualizations.
+The impact of hyper-parameters and other analyses are reported in the supplementary materials.
+5.5.1 Impact of each mask
+As argued, the sparse mask is proposed to assist the detector in identifying local adversarial noises in generative-based adversarial data.
 As shown in Table 5, the absence of the saliency detection model $f_{s}$ or the GradCAM model $f_{c}$ has little effect on the overall system.
 This is because the regions masked by $f_{s}$ and $f_{c}$ are somewhat complementary.
-However, without the Sobel operation to make the mask sparse, the detector fails entirely in detecting attacks, except for CDA, because the noise in generative-based adversarial data tends to be imbalanced and localized, whereas the noise in CDA is more global.
-Without the sparsification provided by the Sobel operation, the generated pseudo-noise becomes too dense, preventing the detector from effectively focusing on local noise.
-Time Cost of Inference
-We conducted experiments to calculate the inference time cost for 100 samples on ImageNet100.
-As shown in Table 6, the time cost of our model is slightly higher than that of LID, LiBRe, and SPAD, and it only takes 0.0485 seconds to process one image.
-For actual use, a minor increase in time is perfectly acceptable in exchange for a significant increase in detection performance.
-Figure 3: 2D T-SNE visualizations. (a) ImageNet100 flattened noises of adversarial data and Perturbation Forgery. (b) ImageNet100 features extracted by the model trained with Perturbation Forgery.
-Functionality of Perturbation Forgery
-To validate the functionality of our method, we extract features and flattened noises from adversarial data and Perturbation Forgery and visualize them using 2D t-SNE projection.
-All adversarial data come from unseen attacks.
-As shown in Figure 3(a), by continuously perturbing the adversarial noise distribution of a commonly used attack, Perturbation Forgery forms an open covering of noises from unseen attacks.
-Individual samples float outside of the collective because they are outliers from the noise distribution.
+However, without the Sobel operation to sparsify the mask, the detector fails to detect most attacks, except for CDA.
+This is because generative-based adversarial noise tends to be imbalanced and localized, while CDA’s noise is more global.
+The generated pseudo-adversarial noise without Sobel becomes too dense, preventing the detector from effectively focusing on localized noise.
+5.5.2 Functionality Visualization
+To validate the functionality of our method, we extract features and flattened noise from adversarial data and Perturbation Forgery-generated data, then visualize them using 2D t-SNE projections.
+All adversarial data originate from unseen attacks.
+As shown in Figure 3(a), Perturbation Forgery continuously perturbs the adv-noise distribution of the initial attack, thereby forming an open covering for noises from unseen attacks.
+Individual samples appear as outliers, floating outside the collective noise distribution.
 As shown in Figure 3(b), by learning to separate natural data from those generated by Perturbation Forgery, the trained detector can effectively distinguish between natural and adversarial data.
-Conclusion
+Figure 3: 2D T-SNE visualizations.
+(a) ImageNet100 flattened noises of adversarial data and Perturbation Forgery-generated data.
+(b) ImageNet100 features extracted by the backbone model trained with Perturbation Forgery.
+Table 6: AUROC scores against white-box attacks under $\epsilon=4/255$ on CIFAR-10.
+Attack BIM PGD DIM MIM AA
+AUROC 0.9180 0.9168 0.9062 0.9031 0.9019
+Table 7: Impact of data size on distribution estimation: AUROC score against PGD under $\epsilon=4/255$ on CIFAR-10.
+Data size 100 250 500 750 1000 2000
+AUROC 0.8278 0.9368 0.9866 0.9921 0.9965 0.9948
+5.5.3 Detecting White-box Attack
+To verify the robustness of our approach to white-box attacks, i.e., both the victim model and the detector are available to the attacker, we conduct the experiment on CIFAR-10.
+Following [3], a modified objective of attacks is:
+$\max_{\boldsymbol{\eta}}\mathcal{L}(f(\hat{\boldsymbol{\mathrm{x}}}),y)+\alpha\cdot\mathcal{L}(f_{d}(\hat{\boldsymbol{\mathrm{x}}}),y_{d}),\ s.t.||\boldsymbol{\eta}||\leq\epsilon$ (12)
+where $y$ is the ground-truth label, $y_{d}$ = 0 indicates that $\hat{\boldsymbol{\mathrm{x}}}$ is an adversarial example, and $\alpha=1$ is a constant balancing between the attack of the classifier $f$ and detector $f_{d}$.
+As shown in Table 6, our method achieves a high average AUROC of 0.9092, demonstrating its robustness against white-box attacks.
+This may be attributed to the comprehensive open covering formed by perturbation forgery, which effectively encompasses adversarial data from white-box attacks, making it harder to mislead the trained detector.
+Figure 4: Impact of the initial attacks (AUROC) on CIFAR-10 under $\epsilon=4/255$. Y-axis: initial attack. X-axis: testing attack.
+5.5.4 Impact of the Initial Attack
+Our proposed perturbation forgery requires a commonly used attack to construct the open covering.
+To assess the impact of the initial attack choice, we train detectors using distributions estimated from various initial attacks and test them against other attacks on CIFAR-10.
+As shown in Figure 4, the detectors consistently achieve a high AUROC score above 0.99, indicating that our method does not depend on a specific attack.
+5.5.5 Data size for Distribution Estmation
+To assess the influence of data size for distribution estimation, we train detectors using distributions estimated from various data sizes and test them against PGD on CIFAR-10.
+As shown in Table 7, our method only requires a small amount of data ($\geq 500$ samples) to estimate the adversarial noise distribution so as to train an effective detector.
+Notably, using fewer than 500 samples leads to a significant decrease in detection performance.
+6 Conclusion
 In this paper, we explore and define the proximity relationship between adversarial noise distributions by modeling these distributions as multivariate Gaussian.
-Based on our assumptions and definitions, we establish a perturbation method to craft proximal distributions from a given distribution and demonstrate the existence of an open covering of adversarial noise distributions.
-After that, we propose the core idea of obtaining a detector with strong generalization abilities against all types of adversarial attacks by learning to distinguish the open covering from the distribution of natural data.
+Based on our assumption and definition, we establish a perturbation method to craft proximal distributions from a given distribution and demonstrate the existence of an open covering of adversarial noise distributions.
+After that, we propose the core idea of obtaining a detector with strong generalization abilities against various types of adversarial attacks by training on the open covering of adversarial noise distributions.
 Based on this insight, we heuristically propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production, to train an adversarial detector capable of detecting unseen gradient-based, generative-based, and physical adversarial attacks, agnostic to any specific models.
-The major limitation of our method is that finding the optimal scale parameters for distribution perturbation is challenging.
-We must rely on continuous experimentation to determine these parameters.
-However, even if these parameters are not optimal, our method still maintains consistent detection performance across various attacks on multiple general and facial datasets, with a satisfactory inference time cost.
-Appendix
-Appendix A Proofs
-Theorem Proofs
-Proof of Theorem 1.
-Proof.
-Denoting the distribution of natural sample $\mathbf{x}$ as $P$, we use the 1-Wasserstein distance as the metric $d(\cdot,\cdot)$, to prove the theorem.
-1) We denote the independent adv-noise distribution of $\eta_{i}$ as $Q_{i}\in\mathcal{P}_{a}$ where $\eta_{i}=\hat{\mathbf{x}}_{i}-\mathbf{x}$, and a zero-distribution as $P$ where its mean and variance are zero vectors.
-Then the 1-Wasserstein distance between $\eta_{i}$ and $\eta_{j}$ can be bound by its metric property and dual form:
-$W_{1}\left(Q_{i},Q_{j}\right)\leq$ $W_{1}(Q_{i},P)+W_{1}(Q_{j},P)$
-$\leq$ $\sup_{\|g\|_{L}\leq 1}\mathbb{E}_{\boldsymbol{\eta}_{i}\sim Q_{i}}[g(\boldsymbol{\eta}_{i})]-\mathbb{E}_{\mathbf{x}\sim P}[g(\mathbf{x})]$
-$+\sup_{\|g\|_{L}\leq 1}\mathbb{E}_{\boldsymbol{\eta}_{j}\sim Q_{j}}[g(\boldsymbol{\eta}_{j})]-\mathbb{E}_{\mathbf{x}\sim P}[g(\mathbf{x})]$
-$\leq$ $2\epsilon,$
-where $\|g\|_{L}$ is the Lipschitz constant of function $g$ w.r.t. the norm induced by the metric here.
-2) Consider two random noise $\boldsymbol{\eta}_{i}\sim\mathcal{N}_{i}(\mu_{i},\sigma_{i}^{2}\mathbf{I}_{d})$ and $\boldsymbol{\eta}_{j}\sim\mathcal{N}_{j}(\mu_{j},\sigma_{j}^{2}\mathbf{I}_{d})$.
-While it’s mean and variance satisfied that $\|\mu_{i}-\mu_{j}\|\leq\epsilon$ and $|\sigma_{i}-\sigma_{j}|\leq\epsilon$.
-then the 2-Wasserstein distance between $\eta_{1}$ and $\eta_{2}$ is
-$W_{2}\left(\boldsymbol{\eta}_{i},\boldsymbol{\eta}_{j}\right)^{2}=$ $\left\|\mu_{i}-\mu_{j}\right\|^{2}+\operatorname{trace}(\sigma_{i}^{2}\mathbf{I}_{d}+\sigma_{j}^{2}\mathbf{I}_{d}$
-$-2\left((\sigma_{j}^{2}\mathbf{I}_{d})^{1/2}\sigma_{i}^{2}\mathbf{I}_{d}(\sigma_{j}^{2}\mathbf{I}_{d})^{1/2}\right)^{1/2})$
-$\leq$ $\epsilon_{\mu}^{2}+d\cdot\epsilon_{\sigma}^{2},$
-where $(\sigma_{i}^{2}\mathbf{I}_{d})^{1/2}$ denotes the principal square root of $\sigma_{i}^{2}\mathbf{I}_{d}$. ∎
-Corollary 1.
-1) All noise distributions proximal to the known adversarial distribution $p(\boldsymbol{\eta})$ form an open $\varepsilon$-ball centered on $p(\boldsymbol{\eta})$, denoting as $\mathcal{P}_{\varepsilon}=\{p(\boldsymbol{\eta}_{i})|d(p(\boldsymbol{\eta}_{i}),p(\boldsymbol{\eta}))<\varepsilon)\}$, and also form a metric space $(\mathcal{P}_{\varepsilon},d)$.
-2) The adv-noise distribution set is also located in this ball: $\mathcal{P}_{a}\subset\mathcal{P}_{\varepsilon}$.
-3) Based on the properties of the complete metric space $(\mathcal{P}_{\varepsilon},d)$, for all subset $\mathcal{P}_{i}\subset\mathcal{P}_{\varepsilon}$, there exists a finite open covering of the subset: $\mathcal{P}_{i}\subset\{G_{c}\}_{c\in J}$.
-Proof.
-1) Consider 1-Wasserstein distance as metric $d$, for $\forall Q_{i},Q_{j}\in\mathcal{P}_{\varepsilon}$, we have
-$W_{1}\left(Q_{i}\right)\geq 0.\quad W_{1}\left(Q_{i},Q_{j}\right)=0\ if\ and\ only\ if\ Q_{i}=Q_{j}.$ (13)
-$W_{1}\left(Q_{i},Q_{j}\right)=W_{1}\left(Q_{j},Q_{i}\right).$ (14)
-$W_{1}\left(Q_{i},Q_{j}\right)\leq W_{1}(Q_{i},P)+W_{1}(Q_{j},P).$ (15)
-Therefore, $(\mathcal{P}_{\varepsilon},d)$ is a metric space.
-2) It is obvious according to the definition of $\mathcal{P}_{\varepsilon}$.
-3) For any subset $\mathcal{P}_{i}=\{p_{i1},p_{i2},\dots,p_{in}\}\subset\mathcal{P}_{\varepsilon}$, we can construct several spherical subsets $\{\mathcal{P}_{ij}\}$ with $p_{i1}$ as the center and $|\varepsilon-d(p_{i1},p(\boldsymbol{\eta}))|$ as the radius.
-$\{\mathcal{P}_{ij}\}$ is a open covering of $\mathcal{P}_{i}$.
-Appendix B Experimental Details
-Dataset
-CIFAR10
-CIFAR10 consists of a training set of 50,000 images and a test set of 10,000 images, each with a resolution of 32x32 pixels.
-These images are divided into 10 different classes.
-The classes represent everyday objects such as airplanes, automobiles, birds, cats, deer, dogs, frogs, horses, ships, and trucks.
-ImageNet100
-ImageNet-100 is a subset of ImageNet (Deng et al.
-2009), comprising 100 classes selected from the original ImageNet dataset, each containing a substantial number of high-resolution images, with 1,300 images per class.
-Makeup
-Makeup comprises images of faces with and without makeup.
-These images are collected from various sources to ensure a diverse representation of facial features, makeup styles, and skin tones.
-There are 333 before-makeup images and 302 after-makeup images.
-Table 7: Detection AUC(%) against various gradient-based attacks on CIFAR-10 ($\epsilon=4/255$).
-Detector BIM PGD RFGSM DIM MIM NIM VNIM SNIM AA
-SPAD (Wang et al. 2023a) 99.51 99.38 99.35 99.52 99.66 99.85 99.47 99.63 99.74
-EPSAD (Zhang et al. 2023) 99.99 99.98 99.98 99.99 99.99 99.99 99.99 99.99 99.99
-ours 99.65 99.65 99.68 99.47 99.80 99.80 99.76 99.64 99.88
-Figure 4: The influence of parameters on the detection performance.
-Figure 5: The influence of attack intensity on the detection performance.
-CelebA-HQ
-CelebA-HQ dataset is a high-quality version of the original CelebA dataset.CelebA-HQ consists of 30,000 high-resolution images of celebrity faces, derived from the CelebA dataset through a progressive GAN-based upsampling and quality enhancement process.
-Each image in CelebA-HQ is 1024x1024 pixels.
-LFW
-LFW contains 13,233 labeled images of 5,749 different individuals collected from the internet, each image is labeled with the name of the person pictured.
-LFW images vary widely in terms of lighting, facial expression, pose, and background, closely reflecting real-world conditions.
-The images are provided in a resolution of 250x250 pixels, and the faces are roughly aligned based on the eye coordinates.
-Adversarial Attacks
-Gradient-based attacks
-We use Torchattacks (Kim 2020) to implement adversarial attacks.
-The victim models used for adversarial data generation are ResNet-50 (He et al.
-2016) for ImageNet, WideReseNet-28 for CIFAR10 (Krizhevsky, Hinton et al.
-2009), and ArcFace (Deng et al.
-2019) for face attacks.
-Generative-model-based attacks
-For CDA, we use the pre-trained generator which has been trained on ImageNet with a relativistic loss against ResNet152.
-For TTP, we use the pre-trained-generators trained against ResNet50 for 8 target labels, they are: Grey Owl(24),Goose(99),French Bulldog(245),Hippopotamus(344),Cannon(471),Fire Engine(555),Parachute(701), Snowmobile(802).
-We generate 1000 adversarial examples for each label and randomly select 125 adversarial examples for each label.
-For M3D, its settings are consistent with TTP.
-For Diff-PGD, we use the global attack to craft adversarial examples, target classifier is ResNet50, the diffusion model accelerator is ddim50, the reverse step in SDEdit is 3, and the iteration number of PGD is 10, the step size is 2.
-For Diff-attack, the target classifier is ResNet50, DDIM sample steps are set as 20, and iterations to optimize the adversarial image are set as 30.
-For the methods mentioned above, we all use the authors’ implementation to craft adversarial examples.
-Face attacks
-For face attacks, 1 gradient-based attack TIPIM (Yang et al.
-2021), two GAN-based attacks including AdvMakeup (Yin et al.
-2021) and AMT-GAN (Hu et al.
-2022), and 3 physical attacks including Adv-Sticker (Tong et al.
-2021), Adv-Glasses (Sharif et al.
-2016), and Adv-Mask (Tong et al.
-2021) produce face adversarial samples by attacking the victim face model ArcFace (Deng et al.
-2019).
-Baselines
-Adversarial detection
-For LiBRe [^1] https://github.com/thudzj/ScalableBDL/tree/efficient/exps, we trained a ResNet-50 on ImageNet100 and finetuned it using the authors’ implementation, the parameters were kept consistent with theirs.
-For EPSAD [^2] https://github.com/ZSHsh98/EPS-AD, we use the pre-trained models provided by the authors and follow the instructions by the authors to get the detection results.
-Due to the time-consuming nature of computing EPS, we evaluated 1,000 images for each attack method on ImageNet-100.
-Synthetic image detection
-Since existing adversarial example detection methods do not address adversarial examples based on generative models, we compared our method with various synthetic image detection methods to effectively validate its effectiveness.
-We conducted detection experiments on adversarial examples based on generative models using the following methods: CNN-Detection(Wang et al.
-2020) [^3] https://github.com/peterwang512/CNNDetection, LGrad(Tan et al.
-2023) [^4] https://github.com/peterwang512/CNNDetection, Universal-Detector(Ojha, Li, and Lee 2023) [^5] https://github.com/Yuheng-Li/UniversalFakeDetect?tab=readme-ov-file#weights, and DIRE(Wang et al.
-2023b) [^6] https://github.com/ZhendongWang6/DIRE.
-We conducted experiments using the pre-trained models and authors’ implementation code of the above methods.
-Implementation Details
-Experiments are implemented using 2 Nvidia Deforce RTX 3090 GPUs. For training detectors, we use an Adam optimizer with a learning rate of 2e-4, momentum of 0.9, and weight decay of 5e-6.
-Data pre-processing
-We resize images to $32\times 32$ for CIFAR10, $256\times 256$ for ImageNet100, Makeup and CelebA-HQ, and $112\times 112$ for LFW.
-For ImageNet100, after resizing, we center-crop the images to $224\times 224$.
-After that, all of the images are resized to $256\times 256$ followed by random horizontal flip and normalization.
-Noise distribution
-As related in the main paper, we flatten the generated noise to a vector and estimate its distribution.
-However, for a noise with a shape of $256\times 256\times 3$, the length of the corresponding flattened noise is $196,608$, making it impossible to run the algorithm on a single GPU.
-Therefore, we fix the size of noise to $32\times 32\times 3$, and concatenate several sampled pseudo noises to fit the larger image.
-Figure 6: Natural images and corresponding pseudo-adversarial images generated by Perturbation Forgery.
-Experiment on CIFAR-10
-We conduct experiments to detect gradient-based attacks on CIFAR10.
-As reported in Table 7, our method demonstrates effective detection (achieving an AUC score over 0.99) across a variety of attack methods, validating our assumptions of the noise distribution proximity.
-Impact of Parameters
-We select typical attacks and present curves in Figure 4 to depict the impact of parameters on the detection performance.
-The parameters $\epsilon_{\mu}$ and $\epsilon_{\sigma}$ have little effect on detecting PGD, while $\gamma_{c}$ and $\gamma_{l}$ have significant effects on detecting M3D and Diff-PGD.
-This might be because, regardless of whether $\epsilon_{\mu}$ and $\epsilon_{\sigma}$ are large or small, the accuracy of forming the open coverage remains unaffected as long as the data is sufficient.
-However, too small thresholds lead to adding pseudo-noise to insensitive and low-frequency regions, while too large thresholds result in extremely sparse pseudo-noise that is far from the adversarial noise distribution.
-To explore the impact of perturbation intensity of attacks, we select typical attacks and present the curves in Figure 5, where $\epsilon_{\mathrm{a}}$ denotes the intensity of unseen attacks and $\epsilon_{\mathrm{i}}$ denotes the intensity of the initial attack for Perturbation Forgery.
-We can see that the detector trained with a minor $\epsilon_{\mathrm{i}}$ is able to detect adversarial data generated with a larger $\epsilon_{\mathrm{a}}$.
-When $\epsilon_{\mathrm{i}}$ is larger, adversarial data generated with a minor $\epsilon_{\mathrm{a}}$ is not easy to detect.
-This might be because adversarial noises generated with a higher $\epsilon_{\mathrm{a}}$ tend to be more obvious to detect.
-When $\epsilon_{\mathrm{a}}$ is too small than $\epsilon_{\mathrm{i}}$, the distance between the initial attack noise and the unseen attack noise is too far away, reducing the detection performance.
-Figure 7: 2D T-SNE visualizations. (a) CIFAR-10 flattened noises of adversarial data and Perturbation Forgery. (b) CIFAR-10 features extracted by the model trained with Perturbation Forgery.
-Appendix C Limitations
-The major limitation of our method is that finding the optimal scale parameters for distribution perturbation is challenging.
-We must rely on continuous experimentation to determine these parameters.
-Specifically, we initiate parameters of uniform distribution as $\epsilon_{\mu}=1$ and $\epsilon_{\sigma}=0.001$, and adjust them using a small step factor.
-Finally, we find the experiment results when $\epsilon_{\mu}=3$ and $\epsilon_{\sigma}=0.005$ are close to optimal.
-However, even if these parameters are not optimal, our method still maintains consistent detection performance across various attacks on multiple general and facial datasets, with a satisfactory inference time cost.
-Appendix D More Visualization Results
-Visualization on CIFAR-10
-We extract features and flattened noises from adversarial data and Perturbation Forgery from CIFAR-10 and visualize them using 2D t-SNE projection in Figure 7.
-Visualization Examples
-Some examples of Perturbation Forgery at $\epsilon=4/255$ are shown in Figure 6.
+Extensive experiments confirm the effectiveness of our approach and demonstrate its strong generalization ability against unseen adversarial attacks, with a satisfactory inference time overhead.